### v1

# Testing Quantumness via Photon Statistics for Time-Bin based Quantum Random Number Generators

Abstract
Randomness is one of the essential components in many fields including cryptography and simulations. Several Quantum Random Number Generator (QRNG) models have been proposed to produce quantum random numbers, which, due to the quantum theory, are more secure than their classical counterparts. However, QRNGs can not produce true random numbers without deterministic classical post-processing. If the underlying distribution of the QRNG is close to a uniform distribution, a small amount of post-processing is sufficient to produce good random numbers retaining quantumness. In this work, we address the randomness and quantumness in the random numbers generated by the QRNGs. We consider two models of QRNGs, which ideally produce random numbers following different distributions (exponential and uniform), and show that, in practice, they are following similar distributions. These empirical photon distributions can be used to test the quantumness of a QRNG. In this letter, we suggest the $\chi^{2}$ goodness-of-fit to test quantumness, as it is known to be an effective method to test if sample data follows a known distribution. We derive a relation when the underlying sampling distributions of the QRNGs will be $\epsilon$-random. Depending on this relation, a suitable post-processing algorithm can be chosen.
Photon Statistics, Randomness, Quantumness, QRNG
Introduction. Random numbers play a crucial role in numerous applications, including computational methods like Monte Carlo simulations and programming Carlson et al. (2015); Metropolis and Ulam (1949), testing principles of physics and chemistry Bell (1964), as well as in the expansive realm of cryptography for creating cryptographic codes or concealing messages Bennett et al. (1992); Shannon (1949), extending to commercial uses such as lottery games and slot machines. In cryptography, randomness is a key input to ensure security in communication. Classical pseudo-random numbers cannot guarantee communication security due to their deterministic generation process. On the other hand, the randomness of the random numbers generated by Quantum Random Number Generators (QRNG) relies on the theory of quantum physics Ma et al. (2016). Different technologies, including shot noise, tunneling, fluctuation, photon arrival time, etc. have been proposed for QRNGs. Many manufacturers are using these technologies to produce their QRNGs. Herrero-Collantes and Garcia-Escartin (2017); ID Quantique (2001); Qutools (2010); Quintessence Labs (2022); Quside (2023); Quantum CTek (2023); Mars Innovation and Department of Electrical and Computer Engineering & Department of Physics, University of Toronto (2023); Quntum Computing Inc. (2023).
Existing time-bin-based QRNG models produce random numbers from exponential Michael A. Wayne and Kwiat (2009); Wayne and Kwiat (2010) and uniform Nie et al. (2014) distributions if the underlying optical devices perform ideally. But, in practice, they cannot be ideal. Therefore, all of the QRNGs require some classical post-processing using deterministic functions. Although such deterministic classical manipulations increase randomness in the generated number, they also reduce quantumness. The amount of classical post-processing should be minimal to retain quantumness in the generated number. For this, the raw number (before classical post-processing) generated by QRNG should be very close to true random.
In this letter, we aim to address photon statistics for practical scenarios, so that we can estimate randomness of the generated raw number. Heavy post-processing should be applied to produce truly random numbers if these raw numbers are less random. Therefore, photon statistics for QRNGs, which we will discuss in this letter, can be considered as one measure of quantumness for photon arrival time-based QRNGs. Depending on this measure, a post-processing algorithm can be chosen to retain the maximum amount of quantumness. Alléaume et al. Alléaume et al. (2004) characterized photon statistics of single-photon sources using the Mandel parameter, and observed that negative values of the Mandel parameter signify non-classical photon statistics for photon sources. Here, we will address the statistics of photons at the detectors. We consider power instability as the error of the source and the attenuators and detection inefficiency as the error of a single photon detector (SPD).
QRNG model. Although producing single photon states is experimentally challenging Imamoḡlu and Yamamoto (1994); Kim et al. (1999), different experimental approaches Brunel et al. (1999); Lounis and Moerner (2000); Santori et al. (2001, 2002); Vučković et al. (2003); Kuhn et al. (2002); Adam and Mechler (2024); Wang and Wang (2024) have been performed to produce single photons. One of the approaches is using a coherent photon source followed by attenuators. This approach produces photon pulses with negligible probability of having multiple photons within a time interval Argillander et al. (2023).
SourceAttenuatorSPDFPGARandom Number
Figure 1: A basic block diagram of the photon arrival time-based QRNG. Coherent pulses, coming from the source, are sent through attenuators to reduce their power so that the probability of having multiple photons within a predefined time interval (typically less than the dead time of SPD) is negligible Argillander et al. (2023). A SPD detects these photons and the detection time (photon arrival time) has been registered. The FPGA provides random numbers based on the registered time stamps.
A coherent photon source produces coherent states which can be represented as a superposition of Fock states by
$|\psi_{\alpha}\rangle=e^{-\frac{|\alpha|^{2}}{2}}\sum_{n=0}^{\infty}\frac{\alpha^{n}\left(\hat{a}^{\dagger}\right)^{n}}{\sqrt{n!}}|0\rangle,$ (1)
where $\alpha$ is a complex number and $\hat{a}^{\dagger}$ is creation operator. The mean photon number corresponding to this state is given by $|\alpha|^{2}$. This state can be approximated as a single-photon state with high attenuation. The probability of $n$ photons within $T$ time follows Poisson distribution Titulaer and Glauber (1966) and is given by
$P(n,T)=\frac{e^{-\mu T}(\mu T)^{n}}{n!},$ (2)
where $\mu$ is the expected number of photons per unit time. A generic block diagram of the photon arrival time-based QRNG is given in Figure 1.
If $\lambda$ is the wavelength of the photon wave and $n$ photons are present within $T$ time, the corresponding photon power would be $P=\frac{n\hbar c}{\lambda T}=\frac{\mu\hbar c}{\lambda}$, where $\hbar$ is the Planck’s constant and $c$ is the speed of light in vacuum. Let $P_{a}$ be the photon power coming out from the attenuators. Then, the expected number of photons per unit time is given by $\mu=\frac{P_{a}\lambda}{\hbar c}$. This attenuated pulse is detected by the SPD. The time of this detection is registered to produce photon arrival time-based random numbers.
[…]
Figure 2: Time diagram for the arrivals of the photons. There are three different methods to generate random numbers from photon arrival times. (I) Photons are detected in a continuous time reference at time $\{t_{0},t_{1},\dots t_{2(i-1)},t_{2i-1},t_{2i},\dots\}$. The $i$-th bit of the random number will be 00 if $\Delta t_{2i-1}<\Delta t_{2i}$ and $1$ if $\Delta t_{2i-1}>\Delta t_{2i}$. (II) Time is divided into bins with equal lengths. If a photon has been detected in the $i$-th bin then $i$ will be the corresponding random number. A new cycle of bins starts after a fixed sleep time ($t_{s}$) of the SPD. (III) An external time reference of cycle $T$ has been used. Each cycle is divided into $N$ equal time-bins $\tau_{1},\tau_{2},\dots,\tau_{N}$. For each cycle, if the first photon has been detected in $i$-th bin then $i$ will be the corresponding random number. SPD will be in sleep mode for the remaining time of the cycle.
There are different methods to produce random numbers from the registered times. Time diagrams for these methods are shown in Figure 2. In Figure 2(I), if the registered times of photon arrivals are $\{t_{0},t_{1},\dots,t_{2i-2},t_{2i-1},t_{2i},\dots\}$, the non-overlapping consecutive time intervals, i.e., $\Delta t_{2i-1}=t_{2i-1}-t_{2i-2},\Delta t_{2i}=t_{2i}-t_{2i-1}$, are used to generate random numbers as follows. If $\Delta t_{2i-1}<\Delta t_{2i}$, the $i$-th bit of the random number is taken as 00. If $\Delta t_{2i-1}>\Delta t_{2i}$, the $i$-th bit is taken as $1$. But in this process the rate of the generation of random number is slow. The length of the random bit-string is half of the photon count. Comparatively fast processes have been developed in the literature Michael A. Wayne and Kwiat (2009); Wayne and Kwiat (2010); Nie et al. (2014). In Figure 2(II), the time is divided into time-bins, each having equal length. If a photon is detected in the $i$-th time-bin, $i$ is taken as the random number. A new cycle of the bins starts after a fixed sleep time of the detector. Since the photon counting process is a Poisson process, the probability, that the waiting time ($t_{w}$) between the arrivals of two consecutive photons is greater than $t$, is given by $P(t_{w}>t)=P(\text{no photon in }(0,t])=P(0,t)=e^{-\mu t}$. Therefore, the random number generated with this procedure follows exponential distribution Michael A. Wayne and Kwiat (2009); Wayne and Kwiat (2010). To overcome this drawback, a new scheme is proposed in Nie et al. (2014). In this scheme, as depicted in Figure 2(III), an external time reference is used. This external time is divided into cycles of length $T$, which is smaller than the dead time of the SPD. The time $T$ is further divided into $N$ equal subintervals, called the time-bins $\tau_{1},\tau_{2},\dots,\tau_{N}$. In each cycle, if the first photon is detected in the $i$-th time-bin, $i$ would be the corresponding random number. SPD would be in sleep mode for the remaining time of the cycle.
Let $n$ photons be present within $T$ time after the attenuation. Since the photon count follows Poisson distribution (2), the probability, that all the $n$ photons are detected in the $i$-th interval, is given by
$P(i;all|n)$ $=\frac{P(n,|\tau_{i}|)P(0,\frac{i-1}{N}T)P(0,T-\frac{i}{N}T)}{P(n,T)}$
$=\frac{\frac{e^{-\mu|\tau_{i}|}(\mu|\tau_{i}|)^{n}}{n!}e^{-\mu\frac{i-1}{N}T}e^{-\mu(1-\frac{i}{N})T}}{\frac{e^{-\mu T}(\mu T)^{n}}{n!}}$
$=\frac{1}{N^{n}}.$ (3)
Here $|\tau_{i}|$ is the length of $i$-th time-bin, which equals $\frac{T}{N}$. Therefore, if only one photon is present within $T$ time after the attenuation then that photon would be detected in the $i$-th bin with probability $\frac{1}{N}$. Thus the photon arrival time follows random distribution over the time-bins. However, this is the case, when exactly one photon is present in $T$ time, and also when the detection is perfect. Since this is not a practical case, next we discuss the photon statistics for some practical scenarios.
Photon statistics for testing quantumness with ideal devices. We first discuss the actual photon statistics in the practical scenario for external reference time-based QRNG (Figure 2(III)). Here, a random number is generated based on the detection time of the first photon in each time cycle $T$. The probability of detecting the first photon in $i$-th bin, when $n$ photons are present within $T$ time, is given by
$P(i;1st|n)=\frac{\sum_{l=1}^{n}P(0,\frac{(i-1)T}{N})P(l,\frac{T}{N})P(n-l,T-\frac{iT}{N})}{P(n,T)}$
$=\frac{\sum_{l=1}^{n}e^{-\mu T\frac{i-1}{N}}\frac{e^{-\mu T/N}\left(\frac{\mu T}{N}\right)^{l}}{l!}\frac{e^{-\mu T\left(1-\frac{i}{N}\right)}(\mu T)^{n-l}\left(1-\frac{i}{N}\right)^{n-l}}{(n-l)!}}{\frac{e^{-\mu T}(\mu T)^{n}}{n!}}$
$=\sum_{l=1}^{n}\binom{n}{l}\left(\frac{1}{N}\right)^{l}\left(1-\frac{i}{N}\right)^{n-l}$
$=\left(1-\frac{i-1}{N}\right)^{n}-\left(1-\frac{i}{N}\right)^{n}.$ (4)
However, we do not know the number of photons present within the period $T$. Therefore, the probability of detecting the first photon in the $i$-th bin is given by
$P(i;1st)$ $=\sum_{n>=0}P(i;1st|n)P(n,T)$
$=P(i;1st|0)P(0,T)+\sum_{n>0}P(i;1st|n)P(n,T)$
$=0+e^{-\mu T}\sum_{n>0}\frac{\left[\mu T\left(1-\frac{i-1}{N}\right)\right]^{n}}{n!}$
$\qquad-e^{-\mu T}\sum_{n>0}\frac{\left[\mu T\left(1-\frac{i}{N}\right)\right]^{n}}{n!}$
$=e^{-(i-1)\frac{\mu T}{N}}-e^{-i\frac{\mu T}{N}}.$ (5)
Photon statistics for testing quantumness with erroneous devices. Till now, we have considered all of the used optical devices as ideal. But in practice, any error in the source and in the attenuators affects photon power and changes the value of the mean photon number per unit time. Again, if the detector is not $100\%$ efficient, The first photon would be detected in the $i$-th time-bin, if at least one photon has been detected in $i$-th bin, and either (A) no photon has been reached at SPD before the $i$-th bin, or (B) $k$ number of photons have reached before the $i$-th bin, but none of them has been detected. Therefore, the probability of detecting the first photon in the $i$-th bin, when $n$ photons are present within $T$ time, is given by
$P_{d}(i;1st|n)$ $=\frac{\sum_{k=0}^{n-1}P(k,\frac{(i-1)T}{N})(1-d)^{k}\sum_{l=1}^{n-k}\{1-(1-d)^{l}\}P(l,\frac{T}{N})P(n-k-l,T-\frac{iT}{N})}{P(n,T)}$
$=\frac{\sum_{k=0}^{n-1}\frac{e^{-\mu T\frac{i-1}{N}}\left(\mu T\frac{i-1}{N}\right)^{k}}{k!}(1-d)^{k}\sum_{l=1}^{n-k}\{1-(1-d)^{l}\}\frac{e^{-\mu T/N}\left(\frac{\mu T}{N}\right)^{l}}{l!}\frac{e^{-\mu T\left(1-\frac{i}{N}\right)}(\mu T)^{n-k-l}\left(1-\frac{i}{N}\right)^{n-k-l}}{(n-k-l)!}}{\frac{e^{-\mu T}(\mu T)^{n}}{n!}}$
$=n!\sum_{k=0}^{n-1}\frac{\left(\frac{(i-1)(1-d)}{N}\right)^{k}}{k!}\frac{1}{(n-k)!}\sum_{l=1}^{n-k}\binom{n-k}{l}\left[\left(\frac{1}{N}\right)^{l}\left(1-\frac{i}{N}\right)^{n-k-l}-\left(\frac{1-d}{N}\right)^{l}\left(1-\frac{i}{N}\right)^{n-k-l}\right]$
$=\sum_{k=0}^{n-1}\binom{n}{k}\left(\frac{(i-1)(1-d)}{N}\right)^{k}\left[\left(1-\frac{i-1}{N}\right)^{n-k}-\left(1-\frac{i-1+d}{N}\right)^{n-k}\right]$
$=\left(1-(i-1)\frac{d}{N}\right)^{n}-\left(1-i\frac{d}{N}\right)^{n},$ (6)
where $d\in(0,1]$ is the detection efficiency. Since the exact number of photons is unknown, the probability of detecting the first photon in the $i$-th bin becomes
$P_{d}(i;1st)=e^{-(i-1)\frac{\mu Td}{N}}-e^{-i\frac{\mu Td}{N}}=(e^{\frac{\mu Td}{N}}-1)e^{-i\frac{\mu Td}{N}}.$ (7)
The remaining probability $1-\sum_{i=1}^{N}P_{d}(i;1st)=e^{-\mu Td}$ is the probability of detecting no photon in a complete reference cycle. In that case, this reference cycle has no contribution to the random number. Therefore the probability mass function of the distribution corresponding to the detection of the first photon in the $i$-th bin is given by
$f(i)=\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}e^{-i\frac{\mu Td}{N}},\;i=1,2,\dots,N.$ (8)
Figure 3: Plots for the probabilities of detecting the first photon in $256$ time-bins. Probabilities for different values of $\mu Td$ as well as uniform probability distribution have been plotted. $\mu$ is the expected number of photons in unit time, $T$ is the time cycle in the external reference time frame, and $d$ is the detector efficiency. If all the optical components are ideal then the distribution should be given by red dashed line.
Figure 3 shows how $\mu Td$ affects the probability (8). Here we consider $N=256$. From this plot, it is clear that the randomness increases as $\mu Td$ decreases.
Next, let us consider QRNG in Figure 2(II). Here, the probability, that the waiting time ($t_{w}$) between the arrival of two consecutive photons lies in the $i$-th bin, is given by
$P((i-1)t_{l}<t_{w}<it_{l})=P(t_{w}>(i-1)t_{l})-P(t_{w}>it_{l})$
$=e^{-(i-1)\mu t_{l}}-e^{-i\mu t_{l}},$ (9)
where $t_{l}$ is the length of each time-bin. If the detection efficiency is given by $d$, this probability is given by
$P_{d}((i-1)t_{l}<t_{w}<it_{l})=e^{-(i-1)\mu t_{l}d}-e^{-i\mu t_{l}d}.$ (10)
Since in the probability distribution of Equation (8), $\frac{T}{N}$ is the length of the time-bins, the corresponding probability mass function is also given by Equation (8) with $T=Nt_{i}$.
Figure 4: Plot representing the relation between $\mu Td$ and $\epsilon$ such that the probability distribution given by Equation (7) is $\epsilon$-random. Randomness reduces as $\mu Td$ increases. This product can be reduced with the proper choice of external reference cycle $T$, and proper tuning of the source, the attenuators and the detector.
Value of $\bm{\mu Td}$ Number of NIST tests passed
$0.05$ $13$-$14/15$
$0.10$ $11$-$13/15$
$0.15$ $10$-$11/15$
$0.20$ $9$-$10/15$
$0.25$ $7$-$9/15$
$0.30$ $7$-$8/15$
$0.35$ $6$-$8/15$
$0.40$ $6$-$7/15$
Table 1: NIST test result for the data simulated with the distribution given by Equation (8). Three files of size $1$ GB have been used for the test for every $\mu Td$. The data for $\mu Td=0.05$ has passed all but the universal statistical test.
Therefore, these two types of QRNGs perform similarly in the practical scenarios. However, if we can ensure that the source is an actual single-photon source, then the perfect random number can be generated using an external time reference with probability distribution $P_{d}(i)=\frac{d}{N}$ (see Equation (3)). Since the distributions given by Equation (8) are evaluated from the quantum properties of photons, this can be used to test the quantumness in a QRNG. For this testing, one may use the $\chi^{2}$ goodness-of-fit test Dodge (2008) on the data generated by QRNGs as follows. Consider null and alternative hypotheses as
H0: The data follow the distribution (8).
[…]
The test statistic is defined as $\chi^{2}=\sum_{i=1}^{N}\frac{(O_{i}-E_{i})^{2}}{E_{i}}$, where $O_{i}$ denotes the frequency of $i$ in the data, $E_{i}=nf(i)$, $n$ is the size of generated data. Then the null hypothesis will be rejected if $\chi^{2}>\chi^{2}_{N-1,\alpha}$, where $\chi^{2}_{N-1,\alpha}$ is the value given in the chi-square table for significance level $\alpha$. Note that, the typical values of the parameters $\mu,T,d,N$ in $f(i)$ are known to the QRNG designer.
We call a probability distribution $P(\cdot)$, with support $S$, as $\epsilon$-random if and only if $|P(i)-P(j)|<\epsilon$, for all $i,j\in S$. Since the distribution (7) is a monotone decreasing function, it becomes $\epsilon$-random if and only if
$P_{d}(1;1st)-P_{d}(N;1st)<\epsilon$
$\implies e^{-\frac{\mu Td}{N}}+e^{-(N-1)\frac{\mu Td}{N}}-e^{-\mu Td}>1-\epsilon.$ (11)
Figure 4 is a visual representation of the relation (11). It clearly shows that the value of $\epsilon$ increases with $\mu Td$. That means, to generate numbers with $\epsilon$-randomness, one needs to carefully choose the external reference cycle $T$, and configure the source, the attenuators and the detector so that relation (11) has been satisfied.
We have generated random numbers simulating the probability distribution given by Equation (8) for different $\mu Td$. Three files of size $1$ GB have been generated for every $\mu Td$. Then we perform the NIST tests over the generated numbers. $100$ bit streams of length $100,000$ have been considered for the test. There are a total of $15$ tests in the NIST testing suite. A perfectly random number should pass all of them. In Table 1, We tabulate the number of tests our simulated data passed. The data for $\mu Td=0.05$ has passed all but the universal statistical test. This table indicates that lower $\mu Td$ values give rise to random numbers that require less post-processing.
Conclusion. In this letter, we discuss photon statistics for time-bin-based quantum random number generation for practical scenarios. We consider two photon arrival time-based QRNG models, one using external time reference and another without such external reference. We show that, although for ideal devices these two models follow different distributions (one of which is perfectly random), these two QRNGs follow a similar distribution when the photon source is not an actual single photon source. We also suggest a test for quantumness using the $\chi^{2}$ goodness-of-fit test. If the data passes the test, one can say that the data has been generated using the quantum properties. We derive a relation when the photon statistics becomes $\epsilon$-random. We found that by making the product of the expected number of photons ($\mu T$) within an external reference cycle ($T$) and the detector efficiency ($d$) small enough, the randomness of the generated raw random number can be increased. Therefore it would require less post-processing, and it would retain maximum quantumness. In practical cases, the dark count of a detector also contributes to the generation of random numbers. In our discussion, we did not consider the dark count. This work can be extended to see the effect of dark counts on the photon statistics. One may note that our statistics are not meant for black-box testing for commercially available QRNGs. Rather, the physicists and engineers may use these statistics in the QRNG design phase before post-processing, for evaluation, fine-tuning and performance optimization.
Acknowledgement. S. Das would like to acknowledge the support received by the Quantum Delta NL KAT-2 project.