### v5

# Evaluating Quantumness, Efficiency and Cost of Quantum Random Number Generators via Photon Statistics

Abstract
This work presents two significant contributions from the perspectives of quantum random number generator (QRNG) manufacturers and users. For manufacturers, the conventional method of assessing the quantumness of single-photon-based QRNGs through mean and variance comparisons of photon counts is statistically unreliable due to finite sample sizes. Given the sub-Poissonian statistics of single photons, confirming the underlying distribution is crucial for validating a QRNG’s quantumness. We propose a more efficient two-fold statistical approach to ensure the quantumness of optical sources with the desired confidence level. Additionally, we demonstrate that the output of QRNGs from exponential and uniform distributions exhibit similarity under device noise, deriving corresponding photon statistics and conditions for $\epsilon$-randomness.
From the user’s perspective, the fundamental parameters of a QRNG are quantumness, efficiency (random entropy and random number generation rate), and cost. Our analysis reveals that these parameters depend on three factors, namely, expected photon count per unit time, external reference cycle duration, and detection efficiency. A lower expected photon count enhances entropy but increases cost and decreases the generation rate. A shorter external reference cycle boosts entropy but must exceed a minimum threshold to minimize timing errors, with minor impacts on cost and rate. Lower detection efficiency enhances entropy and lowers cost but reduces the generation rate. Finally, to validate our results, we perform statistical tests like NIST, Dieharder, AIS-31, ENT etc. over the data simulated with different values of the above parameters. Our findings can empower manufacturers to customize QRNGs to meet user needs effectively.
I Introduction
Random numbers have important roles in numerous applications, including computational methods like Monte Carlo simulations and programming RevModPhys.87.1067 ; MonteCarlo , testing principles of physics and chemistry PhysicsPhysiqueFizika.1.195 , as well as in the expansive realm of cryptography for creating cryptographic codes or concealing messages Bennett ; Crypto , extending to commercial uses such as slot machines and lottery games konstantinou2004electronic ; yamanaka2017enumeration . In cryptography, randomness is a key input to ensure security in communication. Classical pseudo-random number generators (PRNG) cannot guarantee communication security due to their deterministic generation process. On the contrary, the randomness of a true random number generator (TRNG) is rooted in some physical entropy source Schindler2009 ; Stipčević2014 , which may be classical or quantum.
The randomness of a classical true random number generator (CTRNG) depends on the unpredictability of the physical process that generates the entropy. This unpredictability arises from our limited knowledge of the system’s behavior Stipčević2014 ; mannalatha2023comprehensive . However, if future research resolves these limitations, the process could become predictable and would no longer function as a true random number generator. In contrast, a quantum true random number generator (QTRNG, or more commonly, QRNG) exploits the properties of quantum particles and the principles of quantum mechanics ma2016quantum to produce random numbers. According to quantum theory, these numbers are fundamentally random, making them inherently unpredictable. Thus it would remain unpredictable, irrespective of any technological advancement.
Different technologies, including photon arrival time 10.1063/1.2720728 ; 10.1063/1.2961000 ; WayneRNG ; WayneQRNG ; 10.1063/1.3578456 ; 10.1063/1.4863224 ; 7270265 ; zhang2021simple , tunneling bernardo2017extracting ; 10.1063/5.0055955 , fluctuation gabriel2010generator ; PhysRevApplied.3.054004 ; Guo:19 ; ZQVR ; 10.1063/1.5078547 ; gehring2021homodyne , phase noise Qi:10 ; PhysRevE.81.051137 ; PhysRevA.81.063814 ; Xu:12 ; 10.1063/1.4922417 ; Yang:16 ; app10072431 , etc. have been proposed to produce QRNGs RevModPhys.89.015004 ; mannalatha2023comprehensive . Several manufacturers such as ID Quantique quantis , Qutools quRNG , QNuLabs TROPOS , QuintessenceLabs qStream , Mars Innovation quantoss , Quantum Computing Inc. uQRNG , Quside quside , QuantumCTek QCTek are producing QRNGs following these technologies. These QRNGs are commercially available as off-the-shelf black-box devices. As the photon arrival-time-based QRNGs are the most general and most of companies follow this approach, in this work we focus on the analysis of this type of QRNGs only.
Since no black-box randomness-testing algorithm can distinguish between a PRNG and a QRNG, therefore the first step in validating the quantumness [^1] Sub-Poissonian distribution ($mean>variance$) of photon statistics (more details in Section II). of a random number generator (RNG) is to collect information regarding intermediate attenuator and detector parameters and photon statistics. Then it may be possible to devise a strategy to decide whether the device performs in the classical or quantum regime.
In most of the analyses Jones:23 ; Perina:24 ; Perina24 ; PhysRevA.109.043505 of the photon arrival-time-based QRNGs, the photon statistics have been discussed either by comparing the mean and variance of the sample distribution or using the negativity of the Mandel parameter Mandel:79 . For example, Alléaume et al. Alléaume_2004 discussed photon statistics of photon sources using the Mandel parameter given by (Mandel:79, , Eq. (11a)). They calculated values of the Mandel parameter from observed photon count timestamps and concluded that negative values of the Mandel parameter, which indicates sub-Poissonian statistics, signify non-classical photon statistics for photon sources.
Although the above strategy is analytically correct, a negative value of the Mandel parameter for experimental data may not always indicate sub-Poissonian statistics. For example, a finite number of photon count timestamps from coherent light that actually follows Poissonian distribution may result in a negative value of the Mandel parameter. This happens because the mean and the variance of a sample with a finite size may not always be the same (even under approximation) as the mean and the variance of the underlying distribution BERG2005865 . Therefore, calculating the exact value of any parameter from a sample of finite size is not enough to decide the underlying nature of the distribution.
Another approach to test the quantumness of light is to observe photon antibunching RevModPhys.54.1061 experimentally. However, although a stream of antibunched photons indicates the non-classical nature of light, the converse is not necessarily true. In other words, non-classical light can sometimes be observed as bunched photons PhysRevA.41.475 ; fox2006quantum . We explain this in more detail inB. This approach also consists of equality criteria similar to the above photon statistics case. Therefore this is also not enough to decide photon antibunching.
On the other hand, a single-photon source ensures the quantumness of a QRNG RevModPhys.54.1061 ; fox2006quantum . However, generating single photons are experimentally challenging task PhysRevLett.72.210 ; kim1999single . Various experimental methods PhysRevLett.83.2722 ; lounis2000single ; PhysRevLett.86.1502 ; santori2002indistinguishable ; 10.1063/1.1577828 ; PhysRevLett.89.067901 ; adam2024single ; wang2024multimode have been explored to produce single photons. Existing time-bin-based QRNG models utilize coherent photon sources in conjunction with attenuators to produce random numbers from exponential WayneQRNG ; WayneRNG ; 10.1063/1.3578456 or uniform 10.1063/1.4863224 distributions if the underlying optical devices perform ideally and the source generates perfectly single photon. But, in practice, this is not the case. Nie et al. 10.1063/1.4863224 shows that, for a photon arrival-time-based QRNG, the uniformity and hence the min-entropy depends on device parameters such as expected photon number within a fixed time interval as well as on the detection efficiency. Therefore, it is important to explore photon statistics in practical scenarios in more detail.
The QRNG output theoretically follows a perfectly uniform distribution. However, due to the imperfections of the underlying devices, the raw QRNG data may deviate from uniform randomness and therefore require classical post-processing. One may note that deterministic classical post-processing does not increase the randomness of the QRNG output, it only compresses the raw data with an aim to extracting the min-entropy present in it at the cost of the randomness generation rate.
I.1 Motivation
Given that both PRNGs and QRNGs produce sequences of 0’s and 1’s, merely analyzing these sequences is not sufficient to definitively differentiate between the two PhysRevA.109.022243 . Statistical tests like NIST 10.5555/2206233 , Dieharder Die , AIS-$31$ killmann2011proposal , ENT ENT etc. are available to test the randomness of an RNG. However, from the results of these tests, it is not possible to say whether an RNG is quantum or not. So a pertinent question arises:
Q1: How do we test whether an alleged QRNG is truly quantum or not?
Also, there are several schemes of device-independent QRNG (DI-QRNG) pironio2010random ; 10.1145/2885493 ; acin2016certified ; bierhorst2018experimentally ; liu2018device ; PhysRevLett.120.010503 ; 9153851 and certified randomness pironio2010random ; 6326426 ; Passaro_2015 ; acin2016certified ; Nguyen:18 ; 10.1093/oso/9780198788416.003.0008 ; PhysRevX.10.041048 ; zhang2021simple ; Leone:23 ; PhysRevLett.133.020802 ; PRXQuantum.5.020348 based on Bell violation, non-separability, etc. acin2016certified ; Alicki_2008 ; Alicki_2008a ; Martin_2023 . Though they do not rely on trusted device assumptions and do not require any post-processing, the required resources (that is, entangled qubits) are too costly. More recent works Mironowicz_2016 ; 10.1145/3564246.3585145 attempt locally certified randomness without entanglement, but they require random seeds and some kind of bootstrapping that may be considered analogous to classical post-processing.
On the other hand, the optical trusted device QRNG mannalatha2023comprehensive relies on the assumption that the devices used for extracting randomness and the source are well-characterized and no one can change any configuration. This means that no third party can access the QRNG or influence its results. In contrast, single-photon-based QRNGs can still be verified using photon-count statistics PhysRevA.41.475 ; fox2006quantum ; RevModPhys.68.127 ; qute.202100062 improving the security. A detailed comparison between trusted QRNG and DI-QRNG is given in A. In this work, we focus on the QRNG of the optical trusted device type along with a two-fold statistical method to verify its quantumness.
Side-channel attacks 8924447 ; 8942152 ; chowdhury2021physical can be performed in the post-processing step to gather information about the random number. Therefore, to reduce the compression amount of raw data and to prevent the data from side-channel attacks, the amount of post-processing should be minimal to ensure the maximum entropy in raw data with high speed and reduce the possibility of side-channel attacks on post-processing. To achieve this, the raw output number (before post-processing) generated by QRNG should be as close to uniform random as possible. Nie et al. 10.1063/1.4863224 provided a lower bound for the min-entropy of the raw data in the presence of device errors. Depending on this min-entropy, a suitable post-processing algorithm can be chosen to retain the entropy from the quantum source. As all the practical devices are prone to error, this raises an important question:
Q2: How to increase the min-entropy of the raw data in presence of the device errors?
Q1 and Q2 regarding quantumness and randomness are crucial from the manufacturer’s perspective. Besides these, the random number generation rate and the cost of the QRNG are also important parameters from the user’s perspective. To achieve a comprehensive treatise, we analyze QRNG from both perspectives. Depending on the users’ requirements, the trade-off between the randomness of the quantum source, the random number generation rate, and the cost of the QRNG can vary. For example, in the case of secure communication, the randomness from the quantum source should have the highest priority, and also the generation rate should be high. On the other hand, in token generation or lottery games, a low rate of random number generation would be sufficient. Therefore, to fulfill the user’s requirement, the answers to the following questions are required.
Q3: How to choose the optical parameters (a) the photon numbers in unit time, (b) the time cycle and (c) the detection efficiency of a QRNG depending on the user’s requirements?
I.2 Our Contributions
In this paper, we aim to evaluate the quantumness of photon arrival-time-based QRNG from the manufacturer’s perspective as answers to Q1 and Q2 and randomness, random number generation rate, and cost from the user’s perspective as the answer to Q3. A similar analysis can also be performed for other QRNG models. For the manufacturer, we first address the validation of quantum sources using photon counting statistics. Considering that the sample data for the number of photons within a fixed time is available, we suggest a two-fold statistical method to verify whether the source of the sample is quantum or not, overcoming the limitations of the existing approaches discussed above. The method consists of interval estimation and hypothesis testing, to ensure sub- or super-Poissonian distribution up to a desired confidence level. Several photon-number-resolving detectors have been proposed PhysRevA.71.033819 ; Divochiy2008 ; Kardyna2008 ; PhysRevA.99.043822 ; Natarajan2012 ; Schuck2013 ; Hadfield2009 ; 9844820 ; Provaznik:20 ; Marsili2009 and demonstrated 1263782 ; PhysRevA.67.061801 ; 10.1063/1.1596723 ; PhysRevA.71.061803 ; PhysRevA.75.062325 ; Cheng2023 ; Lita:08 ; Calkins:13 ; Reddy:20 ; Marsili2013 ; 10.1063/1.5000001 ; Zhang2017 ; Pernice2012 ; doi:10.1021/acsphotonics.4c01680 ; 10.1063/1.1388868 in recent years that can be used to perform these tests.
As another contribution, also from the manufacturer’s perspective, we derive the photon statistics for two-photon arrival-time-based QRNGs in practical scenarios. For the practical scenario, we consider power instability as the error of the source and the attenuators and detection inefficiency as the error of a single photon detector (SPD).
From the user’s perspective, we see that the maximum entropy from the quantum source may not be achieved due to some practical constraints related to the random number generation rate, the timing error in registering the photon detection time and the cost of the QRNG. In particular, we show the following.
  • If the expected number of photons within a unit of time is small, the min-entropy from the quantum source would be high. However, it would increase the cost of the QRNG, and it would also reduce the random number generation rate.
  • Although a small external reference cycle of the QRNG increases the min-entropy from the quantum source, to minimize the timing error, it should be higher than a specific value. This reference cycle does not have much effect on the cost or random number generation rate of the QRNG.
  • A detector with less efficiency increases the min-entropy from the quantum source and also reduces the cost of the QRNG. However, it also reduces the random number generation rate.
Finally, we simulated raw data depending on photon statistics we have derived and performed statistical tests like NIST 10.5555/2206233 , Dieharder Die , AIS-$31$ killmann2011proposal and ENT ENT on those raw data. We show that these statistical tests support our theoretical deduction regarding randomness from quantum sources. Nie et al. 10.1063/1.4863224 performed NIST tests after post-processing and showed that the post-processed data passed all of the 15 tests from the NIST suite.
This paper is structured as follows. In Section II we discuss the quantumness of a single-photon source. In Section III and IV we explore single-photon-based QRNGs and coherent pulse-based QRNGs respectively and mention their weakness in present days. We also discuss noisy beam-splitters and detectors in the case of single-photon-based QRNG in Section III. Section V reveals our two-fold proposal to check the quantumness of a photon-based quantum device as an answer to Q1. Detailed discussion on coherent pulse-based QRNGs has been performed in Section VI to answer Q2. In this section, we have also mentioned how our findings are different from the work of Nie et al. 10.1063/1.4863224 . Section VII answers Q3 by separately analyzing the effects of the parameters mentioned in that question. The answer to Q3 by combining the randomness with random number generation rate and entropy for maximum rate is presented in Section VIII. Results from different randomness testing suites for our simulated random numbers based on the analysis of Section VI are given in Section IX. Finally, in Section X we conclude our work.
II Quantumness of a Single-photon Source
In this section, we describe what we mean by quantumness in the case of a single-photon source. We first discuss photon statistics and then revisit the existing procedure to distinguish quantum light from classical light using such photon statistics. Here, we validate the quantumness of a photon source by its statistical closeness to an ideal single-photon source. Another way of defining and validating quantumness is through Bell test, quantum discord, non-separability, etc. acin2016certified ; Alicki_2008 ; Alicki_2008a ; Martin_2023 for QRNGs that are based on entangled photon source. As our point of interest is the single-photon source which does not have any entanglement, we are not concerned with these tests in our current work.
Photon detection and counting is an important part of QRNG which is further used in other practical applications such as quantum cryptography and quantum communication protocols Bennett ; Crypto . In general, photon-based QRNGs generate random numbers by detecting single photons. Therefore, photon-counting statistics play an important role in determining whether or not light falls in the quantum regime.
The distribution of the single-photon-counting statistics can be divided into three classes fox2006quantum :
  1. (i)
    sub-Poissonian, i.e., the distribution variance is less than the mean,
  2. (ii)
    Poissonian, i.e., the distribution variance is the same as the mean, and
  3. (iii)
    super-Poissonian, i.e., the distribution variance is greater than the mean.
Classical wave theory of light follows Poissonian and super-Poissonian statistics fox2006quantum . On the other hand, the sub-Poissonian nature of a photon source cannot be explained by the classical wave theory, making it a clear indication of its quantum nature RevModPhys.68.127 ; qute.202100062 . A detailed discussion is given in B.
Let us consider a probability distribution given by,
$P(n)=\begin{cases}\frac{e^{-\alpha\mu}(\alpha\mu)^{\alpha n}}{(\alpha n)!},&\text{if }\alpha n\text{ is integer,}\\ 0,&\text{otherwise}\end{cases}$ (1)
for $n=0,1,2,\dots$ and some positive real numbers $\mu,\alpha$. Then the mean would be $\mu$ and variance would be $\frac{\mu}{\alpha}$. Therefore, this denotes a sub-Poissonian, Poissonian, or super-Poissonian distribution for $\alpha>1,\alpha=1,\alpha<1$ respectively.
Thermal light follows Bose–Einstein distribution fox2006quantum , given by
$P(n)=\frac{1}{\mu+1}\left(\frac{\mu}{\mu+1}\right)^{n}.$ (2)
The mean of this distribution is given by $\mu$, and the variance is $\mu+\mu^{2}>\mu$. Therefore, this is a super-Poissonian distribution and hence thermal light is an example of classical light.
Next consider a perfect single-photon source that emits photons maintaining equal time interval, $\Delta t$ between two consecutive photons. Then the photon count within time $T$ would be the integer $\mu=\lfloor\frac{T}{\Delta t}\rfloor$. In this case, the mean photon count is given by $\mu$, but the variance is $0<\mu$. Therefore, this is an example of non-classical or quantum light.
A coherent photon source produces coherent states which can be represented as a superposition of Fock states by
$|\psi_{\alpha}\rangle=e^{-\frac{|\alpha|^{2}}{2}}\sum_{n=0}^{\infty}\frac{\alpha^{n}\left(\hat{a}^{\dagger}\right)^{n}}{\sqrt{n!}}|0\rangle,$ (3)
where $\alpha$ is a complex number and $\hat{a}^{\dagger}$ is the creation operator. The mean photon number corresponding to this state is given by $|\alpha|^{2}$. This state can be approximated as a single-photon state with high attenuation. The probability of $n$ photons within $T$ time follows Poisson distribution PhysRev.145.1041 given by
$P(n,T)=\frac{e^{-\mu T}(\mu T)^{n}}{n!},$ (4)
where $\mu$ is the mean number of photons per unit time. It can be shown that for this distribution, the variance of the number of photons per unit time is also $\mu$.
Antibunching property can also be used to distinguish quantum light from classical light. Detailed discussion about this is included in B.
III QRNG based on Single-photon Source
An ideal QRNG, based on single photons, uses $/50$ 50 beam splitter jennewein2000fast , where an $R/T$ beam-splitter denotes an optical device that splits an incoming light beam into $R\%$ reflected and $T\%$ transmitted light beams. A standard technique is to place a $/50$ 50 Beam-splitter making an angle of $45^{\circ}$ with the path of the photons. Thus, a single photon either passes through the beam-splitter, or it is reflected with probability $\frac{1}{2}$. Two photon detectors say $D_{0}$ and $D_{1}$, can be placed over two output paths. If a photon is detected by $D_{i}$, $i$ is registered as a random bit. A block diagram of this process is given in Fig. 1.
[…]
Figure 1: A generic block diagram of the single-photon-based QRNG. Ideally, single-photon, coming from the source, are sent through a $/50$ 50 beam-splitter. Two SPDs ($D_{0},D_{1}$) are placed in two paths to detect the photon. Photon may take either path with probability $0.5$. The random bit would be $i$ if the detection happens at $D_{i}$. In practice, beam splitters and detectors cannot be error-free. However, for an $R/T$ beam-splitter, by choosing detectors as in Eq. (7), uniform random numbers can be generated.
III.1 Effect of Imperfect Devices
If the photon source for this QRNG is not a perfect single-photon source, it may happen that at the same time some photons are transmitted through the beam-splitter and some are reflected. Then, both the detectors may detect photons simultaneously, and the random number cannot be generated. To ensure whether the source is a single-photon source or not, one may perform the first step of the two-fold statistical method mentioned in Section V. If this method reveals that the distribution is not sub-Poissonian, then the source cannot be a single-photon source, and that source cannot be used in this QRNG.
Each bit of the random number generated with the above single-photon-based QRNG has probability $\frac{1}{2}$ of being 00 or $1$ when all the underlying hardware components such as source, beam-splitter, SPD, are ideal. So, ideally, it produces uniform random numbers. However, any defect in the beam-splitter or the detectors may change this probability. Let us consider an $R/T$ beam-splitter and the detectors $D_{i}$ with detection efficiency $d_{i}\in(0,1]$. Suppose, $D_{0}$ is placed on the path of the transmitted light, and $D_{1}$ is placed on the path of the reflected light. Then the probability of 00 in the generated random number, that is, the probability of detecting a photon by detector $D_{0}$ would be
$P(D_{0})=\frac{d_{0}}{d_{0}+d_{1}}\frac{T}{100}.$ (5)
Similarly, the probability of $1$ would be
$P(D_{1})=\frac{d_{1}}{d_{0}+d_{1}}\frac{R}{100}.$ (6)
Therefore, to generate random numbers, we require,
$P(D_{0})=P(D_{1})\implies d_{0}T=d_{1}R.$ (7)
If $d_{0}T\neq d_{1}R$, the generated random number would not be uniformly random. To test this randomness, statistical tests like NIST, Dieharder, AIS-$31$, ENT etc. may be used. See Section IX for a discussion on these.
IV Existing QRNG Models based on Coherent Pulse
It is well known that producing single photon states is experimentally challenging PhysRevLett.72.210 ; kim1999single . However, different experimental approaches PhysRevLett.83.2722 ; lounis2000single ; PhysRevLett.86.1502 ; santori2002indistinguishable ; 10.1063/1.1577828 ; PhysRevLett.89.067901 ; adam2024single ; wang2024multimode have been performed to produce single photons. One of the approaches is using a coherent photon source followed by attenuators. This approach produces photon pulses with negligible probability of having multiple photons within a time interval argillander2023quantum . In this process, the photons follow Poissonian statistics. A generic block diagram of the photon arrival time-based QRNG using photons from coherent sources is given in Fig. 2.
SourceAttenuatorSPDFPGARandom Numbers
Figure 2: A basic block diagram of the photon arrival time-based QRNG using weak coherent pulses. Coherent pulses, coming from the source, are sent through attenuators to reduce their power so that the probability of having multiple photons within a predefined time interval becomes negligible argillander2023quantum . A SPD detects these photons and the detection time (photon arrival time) has been registered. The Field Programmable Gate Array (FPGA) provides random numbers based on the registered time stamps.
If $\lambda$ is the wavelength of the photon wave and $n$ photons are present within $T$ time, the corresponding photon power would be
$P=\frac{n\hbar c}{\lambda T}=\frac{\mu\hbar c}{\lambda},$ (8)
where $\hbar$ is the Planck’s constant and $c$ is the speed of light in vacuum. Let $P_{a}$ be the photon power coming out from the attenuators. Then, the expected number of photons per unit time is given by
$\mu=\frac{P_{a}\lambda}{\hbar c}.$ (9)
This attenuated pulse is detected by the SPD. The time of this detection is registered to produce photon arrival time-based random numbers.
[…]
Figure 3: Time diagram for the arrivals of the photons. There are three different methods to generate random numbers from photon arrival times. (I) Photons are detected in a continuous time reference at time $\{t_{0},t_{1},\dots t_{2(i-1)},t_{2i-1},t_{2i},\dots\}$. The $i$-th bit of the random number will be 00 if $\Delta t_{2i-1}<\Delta t_{2i}$ and $1$ if $\Delta t_{2i-1}>\Delta t_{2i}$. (II) Time is divided into bins of equal length. If a photon has been detected in the $i$-th bin then $i-1$ will be the corresponding random number. A new cycle of bins starts after a fixed sleep time ($t_{s}$) of the SPD. (III) An external time reference of cycle $T$ has been used. Each cycle is divided into $N$ equal time-bins $\tau_{1},\tau_{2},\dots,\tau_{N}$. For each cycle, if the first photon has been detected in $i$-th bin then $i-1$ will be the corresponding random number. SPD will be in sleep mode for the remaining time of the cycle.
There are different methods to produce random numbers from the registered times. Time diagrams for these methods are shown in Fig. 3. In Fig. 3(I), if the registered times of photon arrivals are $\{t_{0},t_{1},\dots,t_{2i-2},t_{2i-1},t_{2i},\dots\}$, the non-overlapping consecutive time intervals, i.e., $\Delta t_{2i-1}=t_{2i-1}-t_{2i-2},\Delta t_{2i}=t_{2i}-t_{2i-1}$, are used to generate the $i$-th bit of the random numbers as follows.
[…]
But in this process 10.1063/1.2961000 , the random number generation rate is very slow. The length of the random bit-string is half of the photon count. Comparatively faster processes have been developed in the literature 10.1063/1.2720728 ; WayneQRNG ; WayneRNG ; 10.1063/1.3578456 ; 10.1063/1.4863224 . In Fig. 3(II), the time is divided into time-bins, each having equal length. If a photon is detected in the $i$-th time-bin, $i-1$ is taken as the random number. A new cycle of the bins starts after the detector has a fixed sleep time. Since the photon counting process is a Poisson process, the probability, that the waiting time ($t_{w}$) between the arrivals of two consecutive photons is greater than $t$, is given by
$P(t_{w}>t)$ $=P(\text{no photon in }(0,t])$
$=P(0,t)=e^{-\mu t}.$ (11)
Therefore, the random number generated with this procedure follows exponential distribution WayneQRNG ; WayneRNG ; 10.1063/1.3578456 ; 10.1063/1.4863224 . Another QRNG scheme is proposed in 10.1063/1.2720728 , as depicted in Fig. 3(III), using external time reference. This external time is divided into cycles of length $T$, which is smaller than the dead time of the SPD. The time $T$ is further divided into $N$ equal subintervals, called the time-bins $\tau_{1},\tau_{2},\dots,\tau_{N}$. In each cycle, if the first photon is detected in the $i$-th time-bin, $i-1$ would be the corresponding random number. SPD would be in sleep mode for the remaining time of the cycle.
Let $n$ photons be present within $T$ time after the attenuation. Since the photon count follows Poisson distribution (4), the probability, that all the $n$ photons are detected in the $i$-th interval, is given by
$P(i;all|n)$ $=\frac{P(n,|\tau_{i}|)P(0,\frac{i-1}{N}T)P(0,T-\frac{i}{N}T)}{P(n,T)}$
$=\frac{\frac{e^{-\mu|\tau_{i}|}(\mu|\tau_{i}|)^{n}}{n!}e^{-\mu\frac{i-1}{N}T}e^{-\mu(1-\frac{i}{N})T}}{\frac{e^{-\mu T}(\mu T)^{n}}{n!}}$
$=\frac{1}{N^{n}}.$ (12)
Here $|\tau_{i}|$ is the length of $i$-th time-bin, which equals $\frac{T}{N}$. Therefore, if only one photon is present within $T$ time after the attenuation then that photon would be detected in the $i$-th bin with probability $\frac{1}{N}$. Thus the photon arrival time follows random distribution over the time-bins. However, this is the case, when exactly one photon is present in $T$ time, and also when the detection is perfect. Since this is not a practical case, next we discuss the photon statistics for some practical scenarios.
V Our Proposal of Two-fold Statistical Method for Testing Quantumness
In Section II we have discussed that for the Poissonian distribution, the mean of the photon-counting distribution is the same as its variance. However, it is interesting to note that mean $=$ variance does not always imply that the distribution is Poisson. For example, consider the probability distribution
$P(k)=(1-p)^{k-1}p,k=1,2,3,\dots,$ (13)
for $0\leq p\leq 1$. If $p=\frac{1}{2}$, this distribution also has mean = variance $=2$. But, it is not Poisson.
Also, in experimental methods the means and variances are calculated from finite data samples. The fineness of the sample cannot perfectly reflect the mean and variances of the underlying distribution. Therefore, the comparison between the mean and the variance with equality may not reflect the underlying distribution undoubtedly.
This motivates us to propose a two-fold method to test the quantumness of the photon source as an answer to question Q1. First, a statistical test is performed to verify whether the mean is equal to or less than or greater than the variance. Note that any statistical validation of equality is not exact, but rather approximate in practice. Thus, if the above test validates the mean and the variance to be equal, we propose to perform a second test to validate that the distribution is indeed Poisson. Numerous photon-number-resolving detectors have been proposed PhysRevA.71.033819 ; Divochiy2008 ; Kardyna2008 ; PhysRevA.99.043822 ; Natarajan2012 ; Schuck2013 ; Hadfield2009 ; 9844820 ; Provaznik:20 ; Marsili2009 and experimentally demonstrated 1263782 ; PhysRevA.67.061801 ; 10.1063/1.1596723 ; PhysRevA.71.061803 ; PhysRevA.75.062325 ; Cheng2023 ; Lita:08 ; Calkins:13 ; Reddy:20 ; Marsili2013 ; 10.1063/1.5000001 ; Zhang2017 ; Pernice2012 ; doi:10.1021/acsphotonics.4c01680 ; 10.1063/1.1388868 in recent years, providing viable tools for carrying out these tests. The statistical tools have been used in this section is discussed in C.
V.1 First Phase: Interval Estimation
Since Poissonian and super-Poissonian statistics are explainable using classical theory, only sub-Poissonian distribution can guarantee the quantumness of a quantum device. We suggest a procedure to decide whether the underlying distribution is sub-Poissonian or not by estimating the confidence interval for the mean of the underlying distribution as follows.
Suppose, the number of photon counts within a fixed time is given as a sample of size $n$. Then by the central limit theorem, $\frac{\langle X\rangle-m}{\sigma/\sqrt{n}}$ is approximately standard normal for large $n$ hogg2019introduction , where $\langle X\rangle$ is the random variable corresponding to the sample mean $\langle x\rangle$, and $m,\sigma^{2}$ are mean and variance of the actual photon-count distribution. Let the sample variance be given by $S^{2}$. Then
$s^{2}=\frac{n}{n-1}S^{2}$ (14)
is an unbiased and consistent estimate of $\sigma^{2}$ hogg2019introduction . Therefore, by replacing $\sigma$ by $s$, we can say that $Y=\frac{\langle X\rangle-m}{s/\sqrt{n}}$ is approximately standard normal for large $n$.
Let $1-\epsilon$ be the confidence level. Suppose $y_{\epsilon}$ be such that
[…]
Then we have the probability
$P(-y_{\epsilon}<Y<y_{\epsilon})$
$=P(-y_{\epsilon}<\frac{\langle X\rangle-m}{s/\sqrt{n}}<y_{\epsilon})$
$=P(\langle X\rangle-\frac{sy_{\epsilon}}{\sqrt{n}}<m<\langle X\rangle+\frac{sy_{\epsilon}}{\sqrt{n}}).$ (16)
Therefore, combining Eq. (15) and Eq. (16), we have
$P(\langle X\rangle-\frac{sy_{\epsilon}}{\sqrt{n}}<m<\langle X\rangle+\frac{sy_{\epsilon}}{\sqrt{n}})=1-\epsilon.$ (17)
So, we can say that $(\langle x\rangle-\frac{sy_{\epsilon}}{\sqrt{n}},\langle x\rangle+\frac{sy_{\epsilon}}{\sqrt{n}})$ is an approximate confidence interval for the distribution mean $m$ with confidence level $1-\epsilon$.
Now we can decide the photon count distribution as sub-Poissonian, or super-Poissonian, with confidence level $1-\epsilon$, if the estimated variance, $s^{2}\leq\langle x\rangle-\frac{sy_{\epsilon}}{\sqrt{n}}$, or $s^{2}\geq\langle x\rangle+\frac{sy_{\epsilon}}{\sqrt{n}}$, i.e., $\frac{\langle x\rangle}{s^{2}}\geq 1+\frac{y_{\epsilon}}{s\sqrt{n}}$, or $\frac{\langle x\rangle}{s^{2}}\leq 1-\frac{y_{\epsilon}}{s\sqrt{n}}$ respectively, where $y_{\epsilon}$ is given by Eq. (15). That is, for some given $\epsilon$ and $y_{\epsilon}$ given by Eq. (15),
[…]
distribution with confidence level $1-\epsilon$.
Alternatively, one can perform hypothesis testing to determine whether the light is sub-Poissonian, Poissonian, or super-Poissonian.
V.2 Hypothesis Testing as an Alternative to Interval Estimation
To determine the quantumness one has to perform two consecutive hypothesis testing. For the first test,
$H_{0}:\ m=s^{2},\ H_{a}:\ m<s^{2},$ (19)
and for the second test,
$H_{0}:\ m=s^{2},\ H_{a}:\ m>s^{2},$ (20)
where $m$ is the mean of the actual distribution and $s^{2}$ is given by Eq. (14). For both tests, the test statistic would be $z=\frac{\langle x\rangle-s^{2}}{s/\sqrt{n}}$. For the first test,
case 1: if $z\geq y_{\epsilon}$, where $y_{\epsilon}$ is given by Eq. (15), $H_{0}$ is rejected. That implies the underlying photon distribution is sub-Poissonian.
case 2: if $z<y_{\epsilon}$, we cannot reject $H_{0}$. Also, we cannot accept $H_{0}$ due to lack of evidence. In that case, we have to perform second hypothesis testing with the same statistic. After the second test, if $z\geq y_{\epsilon}$, $H_{0}$ is rejected again, and we can say that the underlying distribution is super-Poissonian.
Note that, this hypothesis testing is equivalent to the above interval estimation technique, and both results in the same decision.
V.3 Second Phase: Goodness-of-fit Test
Since the equality of mean and variance may not imply Poisson distribution, to ensure the Poisson distribution which is an indicator of a coherent source of light, we suggest a second test which is a goodness-of-fit test, that is, a hypothesis testing to decide whether a sample data comes from a Poisson distribution or not.
For testing the quantumness in the generated random number, one may use the $\chi^{2}$ goodness-of-fit test hogg2019introduction ; chisq on the data generated by QRNGs as follows. Consider null and alternative hypotheses as
$H_{0}$: The data follow the distribution (4).
$H_{\text{a}}$: The data do not follow the distribution (4).
The test statistic is defined as
$\chi^{2}=\sum_{n=1}^{\infty}\frac{(O_{n}-E_{n})^{2}}{E_{n}},$ (21)
where $O_{n}$ denotes the frequency of $n$ number of photon in the data, $E_{n}$ denotes the frequency given by Poisson distribution (4). This test statistic follows $\chi^{2}$ distribution with $k=N-1$ degrees of freedom if all the parameters are known, and its density function is given by,
$f_{\chi^{2}}(x)=\frac{x^{\frac{k}{2}-1}e^{-\frac{x}{2}}}{2^{\frac{k}{2}}\Gamma(\frac{k}{2})},\quad x\geq 0,$ (22)
[…]
is the well-known Gamma function hogg2019introduction . The cumulative distribution function in this case is given by,
[…]
Figure 4: Probability density function (22) for $\chi^{2}$ distribution. The critical region, $\chi^{2}>\chi^{2}_{k,\alpha}$, for the $\chi^{2}$ test is denoted by stars, and the region, $\chi^{2}>\chi^{2}_{obs}$, corresponding to the $p$-value is denoted by dots. Since here $\chi^{2}_{k,\alpha}>\chi^{2}_{obs}$, the $p$-value becomes larger than $\alpha$ and the region corresponding to the $p$-value contains the critical region. Therefore, in this case, the null hypothesis is accepted.
Figure 4 is a plot of the probability density function of a $\chi^{2}$ distribution. The critical region ($\chi^{2}>\chi^{2}_{k,\alpha}$) for the $\chi^{2}$ test is denoted by stars, and the region corresponding to the $p$-value ($\chi^{2}>\chi^{2}_{obs}$) is denoted by dots. The area of the critical region is given by
$P(\chi^{2}>\chi^{2}_{k,\alpha})=1-F_{\chi^{2}}(\chi^{2}_{k,\alpha})=\alpha.$ (24)
$\chi^{2}$-test being a one-tailed test, $p$-value for this test is given by
$P(\chi^{2}>\chi^{2}_{obs})=1-F_{\chi^{2}}(\chi^{2}_{obs}),$ (25)
where $\chi^{2}_{obs}$ is the observed value of the test statistic.
If the typical values of the parameters $\mu,T$ in Eq. (4) are known, we have $k=N-1$. However, if they are unknown, in that case, $k=N-2$. Therefore, $H_{0}$ is rejected if $\chi^{2}_{obs}>\chi^{2}_{k,\alpha}$, where $\chi^{2}_{k,\alpha}$ is the value given by $F_{\chi^{2}}(\chi^{2}_{k,\alpha})=1-\alpha$. In other words, $H_{0}$ is rejected if $p$-value is less than $\alpha$.
As in most cases, the equality of the mean and the variance denotes Poisson distribution, for practical purposes, the last $\chi^{2}$ test may not be needed. We mention this test to ensure the completeness of the discussion.
Note that if synthetic data is generated via simulation according to the actual photon statistics distribution, our two-fold method might inaccurately classify this data as coming from a sub-Poissonian quantum source. We do not assert that our two-fold method can classify any off-the-shelf black-box random number generator. Rather, we claim that, given the internal parameters and intermediate distributions from a real device, our two-fold method provides a more rigorous validation of whether the optical source operates in the classical or quantum regime compared to merely comparing the exact equality of mean and variance, thus confirming the quantum nature of the generated random numbers.
VI Detailed Photon Counting Statistics of Coherent Pulse-based QRNG
The statistics discussed in Eq. (12) of Section IV consider that all of the $n$ photons present within time $T$ would be detected in the same time-bin, which is a crude assumption for $n>1$. In this section, we perform a more fine-grained analysis of the first photon arrival time for coherent pulse-based QRNG devices. This analysis helps us to investigate the quality of QRNG in terms of several parameters, as elaborated in Section VII.
VI.1 Time-bin Statistics for Poissonian Photon with Perfect Devices
Currently available photon arrival time-based QRNGs quRNG ; TROPOS use coherent sources, which generate photons that follow Poissonian statistics. Here, a random number is generated based on the detection time of the first photon in each time cycle $T$. The probability of detecting the first photon in $i$-th bin, when $n$ photons are present within $T$ time, is given by
$P(i;1st|n)$ $=\frac{\sum_{l=1}^{n}P(0,\frac{(i-1)T}{N})P(l,\frac{T}{N})P(n-l,T-\frac{iT}{N})}{P(n,T)}$
$=\frac{\sum_{l=1}^{n}e^{-\mu T\frac{i-1}{N}}\frac{e^{-\mu T/N}\left(\frac{\mu T}{N}\right)^{l}}{l!}\frac{e^{-\mu T\left(1-\frac{i}{N}\right)}(\mu T)^{n-l}\left(1-\frac{i}{N}\right)^{n-l}}{(n-l)!}}{\frac{e^{-\mu T}(\mu T)^{n}}{n!}}$
$=\sum_{l=1}^{n}\binom{n}{l}\left(\frac{1}{N}\right)^{l}\left(1-\frac{i}{N}\right)^{n-l}$
$=\left(1-\frac{i-1}{N}\right)^{n}-\left(1-\frac{i}{N}\right)^{n}.$ (26)
This was also observed by Nie et al. (10.1063/1.4863224, , Eq. (2)). However, we do not know the number of photons present within the period $T$. Therefore, the probability of detecting the first photon in the $i$-th bin is given by
$P(i;1st)$ $=\sum_{n\geq 0}P(i;1st|n)P(n,T)$
$=P(i;1st|0)P(0,T)+\sum_{n>0}P(i;1st|n)P(n,T)$
$=0+\sum_{n>0}\frac{e^{-\mu T}(\mu T)^{n}}{n!}\left[\left(1-\frac{i-1}{N}\right)^{n}-\left(1-\frac{i}{N}\right)^{n}\right]$
$=e^{-\mu T}\sum_{n>0}\frac{\left[\mu T\left(1-\frac{i-1}{N}\right)\right]^{n}}{n!}-e^{-\mu T}\sum_{n>0}\frac{\left[\mu T\left(1-\frac{i}{N}\right)\right]^{n}}{n!}$
$=e^{-(i-1)\frac{\mu T}{N}}-e^{-i\frac{\mu T}{N}}.$ (27)
VI.2 Time-bin Statistics for Poissonian Photon with Inperfect Devices
Til now, we have considered all of the used optical devices as ideal. But in practice, any error in the source and in the attenuators affects photon power and changes the value of the mean photon number per unit time. Again, if the detector is not $100\%$ efficient, the first photon would be detected in the $i$-th time-bin, if at least one photon has been detected in $i$-th bin, and either (A) no photon has been reached at SPD before the $i$-th bin, or (B) $k$ number of photons have reached before the $i$-th bin, but none of them has been detected. Therefore, the probability of detecting the first photon in the $i$-th bin, when $n$ photons are present within $T$ time, is given by
$P_{d}(i;1st|n)$ $=\frac{\sum\limits_{k=0}^{n-1}P(k,\frac{(i-1)T}{N})(1-d)^{k}\sum\limits_{l=1}^{n-k}\{1-(1-d)^{l}\}P(l,\frac{T}{N})P(n-k-l,T-\frac{iT}{N})}{P(n,T)}$
$=\frac{\sum\limits_{k=0}^{n-1}\frac{e^{-\mu T\frac{i-1}{N}}\left(\mu T\frac{i-1}{N}\right)^{k}}{k!}(1-d)^{k}\sum\limits_{l=1}^{n-k}\{1-(1-d)^{l}\}\frac{e^{-\mu T/N}\left(\frac{\mu T}{N}\right)^{l}}{l!}\frac{e^{-\mu T\left(1-\frac{i}{N}\right)}(\mu T)^{n-k-l}\left(1-\frac{i}{N}\right)^{n-k-l}}{(n-k-l)!}}{\frac{e^{-\mu T}(\mu T)^{n}}{n!}}$
$=n!\sum_{k=0}^{n-1}\frac{\left(\frac{(i-1)(1-d)}{N}\right)^{k}}{k!(n-k)!}\sum_{l=1}^{n-k}\binom{n-k}{l}\left[\left(\frac{1}{N}\right)^{l}\left(1-\frac{i}{N}\right)^{n-k-l}-\left(\frac{1-d}{N}\right)^{l}\left(1-\frac{i}{N}\right)^{n-k-l}\right]$
$=\sum_{k=0}^{n-1}\binom{n}{k}\left(\frac{(i-1)(1-d)}{N}\right)^{k}\left[\left(1-\frac{i-1}{N}\right)^{n-k}-\left(1-\frac{i-1+d}{N}\right)^{n-k}\right]$
$=\left(1-(i-1)\frac{d}{N}\right)^{n}-\left(1-i\frac{d}{N}\right)^{n},$ (28)
where $d\in(0,1]$ is the detection efficiency. Since the exact number of photons is unknown, the probability of detecting the first photon in the $i$-th bin becomes
$P_{d}(i;1st)$ $=e^{-(i-1)\frac{\mu Td}{N}}-e^{-i\frac{\mu Td}{N}}$
$=(e^{\frac{\mu Td}{N}}-1)e^{-i\frac{\mu Td}{N}}.$ (29)
The remaining probability
$1-\sum_{i=1}^{N}P_{d}(i;1st)=e^{-\mu Td}$ (30)
is the probability of detecting no photon in a complete reference cycle. In that case, this particular reference cycle has no contribution to the random number. Therefore, the probability mass function of the distribution corresponding to the detection of the first photon in the $i$-th bin is given by
$f(i)=\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}e^{-i\frac{\mu Td}{N}},\;i=1,2,\dots,N.$ (31)
Figure 5: Plots for the probabilities of detecting the first photon in $256$ time-bins given by Eq. (31). Probabilities for different values of $\mu Td$ as well as uniform probability distribution have been plotted. $\mu$ is the expected number of photons in unit time, $T$ is the time cycle in the external reference time frame, and $d$ is the detection efficiency. If all the optical components are ideal then the distribution should be given by red dashed line.
Note that $f(1)$ is the probability $P_{1}$ in Ref. (10.1063/1.4863224, , Eq. (4)). As their aim was to find the minimum entropy for deciding the amount of post-processing, they only derived an upper bound to $P_{1}$. However, we aim to reduce the amount of post-processing. Therefore, we have derived the complete probability mass function to see how the device parameters, $\mu,T$ and $d$ affect the distribution. We also discuss the effect of all the parameters separately in Section VII. Figure 5 shows how $\mu Td$ affects the probability (31). Here we consider $N=256$. From this plot, it is clear that the uniformity increases as $\mu Td$ decreases.
Although this uniformity can be increased by post-processing the raw data, as mentioned above, this reduces the random number generation rate and increases the possibility of side-channel attack. Since here we want to minimize the post-processing, the main goal in designing a QRNG is
$\underset{\mu,T,d}{minimize}\ g(\mu,T,d):=\mu Td,$ (32)
$subject\ to\ \mu>0,T>0,d>0.$
However, this minimization is bounded by some constraints related to the generation rate, the cost and the timing error in registering the photon detection time. These constraints are discussed in Section VII.
Next, let us consider QRNG in Fig. 3(II). Here, the probability, that the waiting time ($t_{w}$) between the arrival of two consecutive photons lies in the $i$-th bin, is given by
$P((i-1)t_{l}<t_{w}<it_{l})$
$=P(t_{w}>(i-1)t_{l})-P(t_{w}>it_{l})$
$=e^{-(i-1)\mu t_{l}}-e^{-i\mu t_{l}},$ (33)
where $t_{l}$ is the length of each time-bin. If the detection efficiency is given by $d$, this probability is given by
$P_{d}((i-1)t_{l}<t_{w}<it_{l})=e^{-(i-1)\mu t_{l}d}-e^{-i\mu t_{l}d}.$ (34)
Since in the probability distribution of Eq. (VI.2), $\frac{T}{N}$ is the length of the time-bins, Eq. (34) is identical with Eq. (VI.2). Therefore, the probability mass function corresponding to the probability Eq. (34) is also given by Eq. (31) with $T=Nt_{i}$.
Therefore, these two types of QRNGs, perform similarly in practical scenarios. However, if we can ensure that the source is an actual single-photon source, then the perfect random number can be generated using an external time reference with probability distribution $P_{d}(i)=\frac{d}{N}$ (Ref. Eq. (12)). Since the distributions given by Eq. (31) are evaluated from the quantum properties of photons, this can be used in the goodness-of-fit test mentioned in Subsection V.3 to test the quantumness in the raw data generated by a QRNG.
Definition 1.
We define a probability distribution $P(\cdot)$, with support $S$, as $\epsilon$-random if and only if
$\max_{V\subseteq S}\left|\sum_{i\in V}\left(P(i)-\frac{1}{|S|}\right)\right|=\frac{1}{2}\sum_{i\in S}\left|P(i)-\frac{1}{|S|}\right|$ $=\epsilon.$ (35)
Figure 6: Plot representing the relation between $\mu Td$ and $\epsilon$ such that the probability distribution is $\epsilon$-random given by Eq. (36). Randomness reduces as $\mu Td$ increases. This product can be reduced with the proper choice of external reference cycle $T$, and proper tuning of the source, the attenuators and the detector.
The distribution (31) becomes $\epsilon$-random if and only if
$\frac{1}{2}\sum_{i\in S}\left|f(i)-\frac{1}{N}\right|=\epsilon.$ (36)
Figure 6 is a visual representation of the Eq. (36) with $N=256$. It clearly shows that the value of $\epsilon$ increases with $\mu Td$. That means, to avoid post-processing while generating random numbers with $\epsilon$-randomness for a given $\epsilon$, one needs to carefully choose the external reference cycle $T$, and configure the source, the attenuators and the detector so that relation (36) has been satisfied.
We have performed the NIST, Dieharder, AIS-31 and ENT tests on the data generated by simulating the probability distribution given by Eq. (31). These test results are provided in Section IX.
VII Effect of $\mu,T,d$ on random number generation rate and cost
As discussed in Eq. (32) of Subsection VI.2, we have to minimize $\mu Td$ to achieve maximum randomness from the quantum source increases. In that discussion, we have considered the product $\mu Td$ as a single quantity. In this section, we separately discuss the effect of
  • the expected photon number in unit time, $\mu$,
  • the external reference cycle, $T$, and
  • the detection efficiency, $d$
on the rate of random number generation, $r$ as well as on the cost of the QRNG.
The randomness of an RNG is characterized by the entropy of the underlying probability distribution followed by the RNG. Let $H(Q)$ be the entropy generated by a quantum source. After postprocessing it becomes
$H(Q)\xrightarrow{postprocessing}H(E),$ (37)
where $H(E)$ is the final entropy generated by the QRNG. As deterministic classical post-processing cannot improve the entropy, in general $H(E)\leq H(Q)$.
Definition 2.
An RNG is said to be ideal if $H(E)=H(U)$, where $H(U)$ is the entropy of uniform randomness. A QRNG is said to be ideal if $H(U)=H(E)=H(Q)$.
However, in the practical scenario, imperfections within used hardware components make ideal QRNG impossible. Therefore, the main goal in designing a QRNG is
[…]
The Shannon entropy 6773024 ; 6773067 of a uniform distribution over $n$ points is given by
$H(U)=-\sum_{i=1}^{n}\frac{1}{n}\log\frac{1}{n}=\log n.$ (39)
On the other hand, the entropy of the probability distribution (31) is given by
$H(Q)=\frac{1-(N+1)e^{-\mu Td}+Ne^{-(N+1)\frac{\mu Td}{N}}}{(1-e^{-\mu Td})(1-e^{-\frac{\mu Td}{N}})}\log\left(e^{\frac{\mu Td}{N}}\right)-\log\left(\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}\right).$ (40)
A detailed derivation of this entropy along with the definition of Shannon entropy and minimum entropy has been added in D. Clearly, $H(Q)\neq\log N=H(U)$. So, we have to apply post-processing as mentioned in Eq. (37) to achieve $H(E)=H(U)$ for some uniform distribution $U$ whose support contains less than $N$ points. This post-processing, also known as randomness extractor, is a data compression procedure using universal hashing PhysRevA.87.062327 . The minimum entropy $H_{min}$ of the raw data decides the amount of this data compression. If $H_{min}\geq\frac{m}{n}$ for some positive integers $m,n$, an $n$-bit raw data needs to be compressed to an $m$-bit data to make it uniform random 10.1063/1.4863224 . As $f(i)$ in Eq. (31) is a decreasing function, the minimum entropy of the photon arrival-time-based QRNG discussed above would be
$H_{min}$ $=-\log f(1)=-\log\frac{1-e^{-\frac{\mu Td}{N}}}{1-e^{-\mu Td}}$
$=-\log\left(1-e^{-\frac{\mu Td}{N}}\right)+\log\left(1-e^{-\mu Td}\right).$ (41)
Nie et al. (10.1063/1.4863224, , Eq. (5)) gave a lower bound of this $H_{min}$. Since our aim is to reduce the amount of post-processing, which depends on $\mu Td$, we want to investigate $\mu,T$ and $d$ separately.
Figure 7: Total entropy (40) and minimum entropy (41) of the probability distribution (31). When the product of the expected number of photons, $\mu T$, within an external reference cycle $T$, and the detection efficiency $d$ is very small, $H(Q)$ and $H_{min}$ are very close to the entropy $H(U)=8$ of the uniform random distribution. Therefore, if $\mu Td$ is very small, the raw data generated by the QRNG would be close to the uniform random number.
Fig. 7 shows how the total entropy (40) and minimum entropy (41) changes with $\mu Td$. Here we have chosen $N=256$. Therefore, the entropy (39) of the uniform distribution would be $H(U)=\log 256=8$. From Fig. 7, it can be easily seen that, if $\mu Td$ is very small, the entropy $H(Q)$ and $H_{min}$ are very close to $8$, which is the entropy of the uniform distribution. In this section, by randomness we will refer to the randomness due to $H_{min}$.
VII.1 Effect of the Expected Photon Numbers
Here we discuss the effect of the expected photon numbers on the cost as well as the random number generation rate of a QRNG.
VII.1.1 Effect of the Exected Photon Number on Cost
Since $\mu$ is the expected number of photons in unit time, this can be reduced by using heavy attenuation. This will directly affect the cost. Heavy attenuation requires a large number of attenuators, increasing the production cost for the random numbers. Thus
small value of μ high production $\text{small value of $\mu\implies$ high production cost}.$ (42)
VII.1.2 Effect of the Expected Photon Number on Random Number Generation Rate
A small value of $\mu$, which indicates a small number of photons, leads to a high chance of no photon detection in a particular external reference cycle of time interval $T$. That means those intervals will have no contribution to the final random number. Therefore, we require more cycles to produce random numbers of expected length. This will reduce the speed of the RNG. From Eq. (VI.2), the probability, $P(0)$, of no photon detection within time $T$ is given by
$P(0)=1-\sum_{i=1}^{N}P_{d}(i;1st)=e^{-\mu Td},$ (43)
where $P_{d}(i;1st)$ is the probability of detecting the first photon in the $i$-th bin.
Suppose, for some $\mu$, the expected number of external cycles to produce a random number of length $k$ bits, where each cycle contains $N$ time-bins, is given by $N_{c}$. Then $N_{c}=\langle X\rangle$, where $X$ is a random variable, whose probability distribution is negative binomial distribution, given by
$P(X=x)=\binom{x-1}{k^{\prime}-1}\left(1-P_{\mu}(0)\right)^{k^{\prime}}P_{\mu}(0)^{x-k^{\prime}},$ (44)
for $x=k^{\prime},k^{\prime}+1,k^{\prime}+2,\dots$, where $k^{\prime}=\lceil\frac{k}{\log N}\rceil$ is the minimum number of cycles required to produce $k$-bit random number and $P_{\mu}(0)=e^{-\mu Td}$. Therefore,
$N_{c}=\langle X\rangle=\frac{k^{\prime}}{1-P(0)}=\frac{k^{\prime}}{1-e^{-\mu Td}}.$ (45)
Thus, the rate of generating random numbers is given by
$r=\frac{k}{TN_{c}}\approx\frac{1-e^{-\mu Td}}{T}\log N\text{ bits per unit time}.$ (46)
Since the external time cycle $T$ is very small, Eq. (46) can be written as
$r\approx\mu d\log N\text{ bits per unit time}.$ (47)
Figure 8: Plot shows how random number generation rate (bit per unit time) changes with the expected photon number per unit time ($\mu$) given by Eq. (47) when $d$ is constant. Five different plots have been shown for five different values of detection efficiency ($d$). From the plot, it can be easily seen that if $\mu$ decreases, the random number generation rate of the QRNG will also decrease.
Clearly, if $\mu$ decreases, this rate will also decrease. The graph in Fig. 8 shows how this rate changes with $\mu$ for fixed $T$ and $d$. Therefore, we cannot reduce $\mu$ arbitrarily to get the maximum amount of randomness.
VII.2 Effect of the timing error of the device on the External Reference Cycle
If we choose the external reference cycle $T$ very small, the product $\mu Td$ would also be small. This should increase the randomness, as discussed in Eq. (32) of the Subsection VI.2. However, a small $T$ would lead to a small length $\frac{T}{N}$ for the time bins. Now, if there are maximum $\delta_{t}$ errors in registering the photon-arrival-time, due to small time-bins, the time may be registered in a time-bin that is different from the actual time-bin as shown in Fig. 9. Let us consider that the actual detection happens in $i$-th bin, that is, within time interval $[(i-1)\frac{T}{N},i\frac{T}{N})$. If the actual detection happens within time interval $[(i-1)\frac{T}{N},(i-1)\frac{T}{N}+\delta_{t})$, the time may be registered as $(i-1)$-th bin. Similarly, if the actual detection happens within time interval $[i\frac{T}{N}-\delta_{t},i\frac{T}{N})$, the time may be registered as $(i+1)$-th bin. However, if the actual detection lies within the time interval $[(i-1)\frac{T}{N}+\delta_{t},i\frac{T}{N}-\delta_{t})$, there will be no error.
[…]
Figure 9: Due to the timing error $\delta_{t}$, the actual may be read as the shifted signal and the registered time-bin may be different from the actual time-bin of photon detection. This would add an error in the generated random number. To minimize such error, we put a lower bound on the external reference cycle $T$ using Eq. (48).
Now let us consider that $T$ is chosen to satisfy the relation $\frac{T}{N}\geq k\delta_{t}$, for some suitable $k$. Then the probability, $p_{e}$, that the detection happens in $i$-th bin but outside of the interval $[(i-1)\frac{T}{N}+\delta_{t},i\frac{T}{N}-\delta_{t})$ is bounded above as $p_{e}\leq\frac{2}{k}$. Here we have assumed uniform distribution as the time $\frac{T}{Nk}$ is very small. Therefore, for a fixed error tolerance, $p_{tol}$, $k$ can be chosen as $k_{tol}=\frac{2}{p_{tol}}$. In that case, the minimum reference cycle $T_{min}$ should be
$T_{min}=Nk_{tol}\delta_{t}.$ (48)
Therefore, we cannot make $T$ arbitrarily small to get the maximum amount of randomness.
VII.3 Effect of the Reference Cycle on Random Number Generation Rate and Cost
Equation (47) indicates that $T$ does not have much effect on the rate of generating random numbers. Also, since $T$ has no impact on the used optical devices, it does not affect cost.
VII.4 Effect of Detection Efficiency
In this subsection, we discuss the effect of the detection efficiency of the detector used in the QRNG on the cost as well as the random number generation rate of a QRNG.
VII.4.1 Effect of the Detection Efficiency on Cost
The value of $\mu Td$ can be decreased by choosing a detector with bad detection efficiency. It is very natural that if a detector can efficiently detect photons, its cost will also be high. Since the randomness increases if $\mu Td$ decreases, choosing small $d$ we can increase the randomness. This will also reduce the cost of the QRNG. Thus
small value of $\text{small value of }d\implies\begin{cases}\text{high randomness}\\ \text{low cost}\end{cases}$ (49)
VII.4.2 Effect of the Detection Efficiency on Random Number Generation Rate
Figure 10: Plot shows how random number generation rate (bit per unit time) changes with the detection efficiency ($d$) given by Eq. (47) when $\mu$ is constant. Five different plots have been shown for five different values of $\mu$, the expected photon number per unit time. From the plot, it can be easily seen that if $d$ decreases, the random number generation rate of the QRNG will also decrease.
Although a bad detector with a small detection efficiency, $d$, reduces production costs, this leads to a high chance of no photon detection in a particular external reference cycle. From Eq. (47) it can be easily seen that if $T$ decreases, the rate of generating random numbers will also decrease. The graph in Fig. 10 shows how this rate changes with $d$ for fixed $T$ and $\mu$. Therefore, it is not possible to choose a detector with arbitrarily small detection efficiency.
VIII Performance and Cost of the QRNG
Figure 11: Plot showing relation between random number generation rate, $r$, expected photon number in unit time, $\mu$, and detection efficiency, $d$ given by Eq. (47). If the values of $\mu$ and $d$ are small, the generation rate is also small. Therefore, to generate random numbers with high rate, we have to choose both $\mu$ and $d$ large.
We can see from Eq. (47) that random number generation rate $r\approx\mu d\log N$, where $\mu$ is the expected number of photons in unit time, $d$ is the efficiency of detector and $N$ is the number of time-bins. Figure 11 is a surface plot of Eq. (36) to show how the random number generation rate of the QRNG changes with $\mu$ and $d$. This shows that we have to consider high values of $\mu$ and $d$ to get a high rate. On the other hand, in Eq. (32) of Subsection VI.2 we show that to get maximum randomness from the quantum source, we have to minimize the product $\mu Td$ small. Also, from Eq. (42)and (49), we can write the cost of the QRNG as
$\text{cost}=\frac{\alpha}{\mu}+\beta d,$ (50)
for some positive constants $\alpha$ and $\beta$. Therefore, this needs some optimization while choosing $\mu$ and $d$.
Considering all the discussions in the previous section, we can now write random number generation rate $r$ as a function of randomness. From Eq. (31), (36) and (47), we have
$2\epsilon$ $=\sum_{i=1}^{N}\left|\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}e^{-i\frac{\mu Td}{N}}-\frac{1}{N}\right|$
$=\sum_{i\leq N/2}\left(\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}e^{-i\frac{\mu Td}{N}}-\frac{1}{N}\right)+\sum_{i>N/2}\left(\frac{1}{N}-\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}e^{-i\frac{\mu Td}{N}}\right)$
$=\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}\left(\sum_{i\leq N/2}e^{-i\frac{\mu Td}{N}}-\sum_{i>N/2}e^{-i\frac{\mu Td}{N}}\right)$
$=\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}\left(\frac{e^{-\frac{\mu Td}{N}}\left(e^{-\frac{\mu Td}{2}}-1\right)}{e^{-\frac{\mu Td}{N}}-1}-\frac{e^{-\frac{\mu Td}{N}\left(\frac{N}{2}+1\right)}\left(e^{-\frac{\mu Td}{2}}-1\right)}{e^{-\frac{\mu Td}{N}}-1}\right)$
$=\frac{\left(1-e^{-\frac{\mu Td}{2}}\right)^{2}}{1-e^{-\mu Td}}=\frac{1-e^{-\frac{\mu Td}{2}}}{1+e^{-\frac{\mu Td}{2}}}.$ (51)
Therefore we can write,
$2\epsilon\left(1+e^{-\frac{\mu Td}{2}}\right)=1-e^{-\frac{\mu Td}{2}}$
$\implies$ $e^{\frac{\mu Td}{2}}=\frac{1+2\epsilon}{1-2\epsilon}$
$\implies$ $r=\frac{2\log N}{T}\ln{\frac{1+2\epsilon}{1-2\epsilon}}.$ (52)
Figure 12: Maximum rate is plotted against $\epsilon$ as Eq. (53) with $T_{min}=Nk_{tol}\delta_{t}=1$ unit of time. It shows that the maximum rate increases very quickly with $\epsilon$. Since from Definition 1, randomness decreases with the increment of $\epsilon$, the maximum rate decreases as randomness increases.
Equation (52) gives the random number generation rate when the generated number is $\epsilon$-random, according to Definition 1. Therefore using Eq. (48), we can write the maximum rate for some given $\epsilon$ as
[…]
where $N$ is the number of time-bins, $\delta_{t}$ is the maximum timing error in photon detection and $k_{tol}$ is a parameter corresponding to the tolerance of this timing error. Fig. 12 plots this maximum rate against $\epsilon$. It shows that the rate decreases with the increment of randomness.
Also, we can write the minimum entropy (41) as a function of $\epsilon$ as
$H_{min}=\log\frac{8\epsilon}{(1+2\epsilon)^{2}\left(1-\left(\frac{1-2\epsilon}{1+2\epsilon}\right)^{\frac{2}{N}}\right)}$ (54)
In fig. 13 we plot this minimum entropy against $\epsilon$.
Figure 13: Minimum entropy, Eq. (54), is plotted against $\epsilon$ for $N=256$. It shows that the minimum entropy decreases, which is a clear indication of decrement of randomness, as $\epsilon$ increases.
Parameters Effects on
$\downarrow$ Randomness Rate Cost
$\mu$uparrow $\downarrow$ $\uparrow$
$T$uparrow -- --
$d$uparrow $\downarrow$ $\downarrow$
Table 1: The effects of decreasing the expected photon count $\mu$, the external reference cycle $T$ and the detection efficiency $d$. A down-arrow, $\downarrow$, denotes the decrement of the corresponding quantity. On the other hand, an up-arrow, $\uparrow$, denotes the increment. $T$ does not affect speed or cost. If the parameters increase, the effects will be in the reverse direction.
Figure 14: Heat map denoting how (a) $\epsilon$-randomness, (b) minimum entropy, (c) random number generation rate and (d) cost varies depending on the expected photon number per unit time, $\mu$, external reference cycle, $T$ and the detection efficiency, $d$. Figure (a) and (b) show the $\epsilon$-randomness in Eq. (51) and minimum entropy in Eq. (41), respectively, indicating that we must keep the product $\mu Td$ small to get high randomness. However, (b) for Eq. (47), with $N=256$, shows if the $\mu d$ is small, the generation rate will be minimal. On the other hand, from (c) for Eq. (50) with $\alpha=1/20,\beta=40$, low-cost demands a small value of $d$ and a high value of $\mu$.
The above discussion has been described as a heat map in Fig. 14. In the figure, dark shade denotes low value and light shade denotes high value of $\epsilon$, minimum entropy, generation rate and cost in Figures 14(a), 14(b), 14(c) and 14(d), respectively. Figure 14(a) and (b) shows that high randomness demands a small value of the product $\mu Td$. However, the product $\mu d$ cannot be very small as Fig. 14(b) indicates. A small $\mu d$ would result in a low random number generation rate. On the other hand, to keep the cost of the QRNG low, we need a small $d$ but a large $\mu$.
In Table 1, we summarize the discussion in tabular form. A down-arrow ($\downarrow$) indicates a decrease in the value of the corresponding quantity, while an up-arrow ($\uparrow$) signifies an increase. The external reference cycle $T$ does not influence speed or cost. If the parameters are increased, the effects will reverse accordingly.
We have already mentioned in Eq. (38) that the main goal in designing a QRNG is to keep $H(Q)$, the entropy from the quantum source as close as $H(U)$, the entropy of uniform randomness. However, we also discussed that this is not the only goal. There are some other constraints as well like random number generation rate and cost. A QRNG should generate random numbers with high rates and low costs. The randomness and the rate together describe the performance of the QRNG. Therefore, the goal in designing a QRNG is to maximize the performance and minimize the cost by choosing appropriate $\mu,T$ and $d$.
IX Results of Randomness Tests
Figure 15: NIST result. $p$-values for passed tests are plotted as bar charts for (a) $\mu Td=0.05$, (b) $\mu Td=0.10$, (c) $\mu Td=0.15$, (d) $\mu Td=0.20$, (e) $\mu Td=0.25$, (f) $\mu Td=0.30$, (g) $\mu Td=0.35$, (h) $\mu Td=0.40$ , (i) $\mu Td=0.45$, (j) $\mu Td=0.50$. For each $\mu Td$, three $1$ GB files are generated by simulating probability distribution (31), and the NIST tests have been performed on those files. Different textures denote results corresponding to different files. A test is considered as passed if the corresponding $p$-value is at least $0.01$. Tests that fail for all of the three files are not included in the bar chart. The vertical dashed line denotes $p$-value $=0.01$
Random numbers and their randomness testing have been discussed in E. The randomness of the random numbers is tested using statistical tests. Here we used four common statistical testing suites namely NIST, Dieharder, AIS-31 and ENT.
Figure 16: NIST test results for the data simulated with the distribution given by Eq. (31). Three files, each containing $1$ GB random number have been used for the test for every $\mu Td$. A test fails if the corresponding $p$-value is found less than $0.01$. The range in the second column arises due to different results from three different files. The number of passed tests decreases with the increment of $\mu Td$. This indicates that as $\mu Td$ increases, the randomness in the generated random number decreases.
The NIST Randomness Testing Suite is an extensive set of statistical tests developed by the National Institute of Standards and Technology (NIST) to evaluate the quality of randomness in random numbers. It is one of the most widely used suites for testing randomness. It contains $15$ different tests to examine different aspects of randomness. Most of the tests have the standard normal or $\chi^{2}$ as reference distribution for their test statistic. Here null hypothesis is taken as the number to be tested is random. For each test, the significance level is considered as $0.01$. If the value of the test statistic for some tests falls in the critical region, or equivalently, calculated $p$-value is less than $0.01$, then the corresponding tests fail rejecting the null hypothesis. A perfectly random number should pass all of the $15$ tests. As mentioned in Section V, the central limit theorem is applicable only for large $n$, tests that use the central limit theorem to determine the reference distribution, have a requirement for the bit stream, that is, the length of the testing sequence, to be large enough (at least $10^{6}$).
The Dieharder tests represent a battery of statistical tests used to assess the quality and randomness of a RNG. This suite contains around $30$ tests for the assessment. These tests consider KS test of uniformity to obtain the $p$-values. Depending on these $p$-values, the decision (passed, weak or failed) has been taken. The results that do not pass and also do not have enough evidence to declare as fail have been mentioned as weak.
AIS-$31$ is a standardized suite of statistical tests for assessing RNGs. This suite comprises a total of 9 tests, typically organized into two main groups. Procedure A, consists of 6 tests aimed at detecting statistically inconspicuous behavior in RNG output. Procedure B, focuses on evaluating the internal random bits of TRNGs through 3 specific tests. These tests ensure that RNGs meet rigorous statistical criteria for randomness, providing a reliable measure of their quality and suitability for applications requiring secure and unpredictable random numbers.
ENT performs $5$ tests on the stream of bytes and produces the outputs for each test. It calculates entropy, $\chi^{2}$ statistic value, the arithmetic mean of the bytes, the Monte Carlo value of $\pi$ and the serial correlation coefficient. Depending on these values user can decide the randomness.
Figure 17: (a) Dieharder and (b) AIS-$31$ test results for the data simulated with the distribution given by Eq. (31). For Dieharder, there are $30$ tests; for AIS-$31$, $9$ tests are there. These tests also reveal that the number of passed tests decreases as $\mu Td$ increases indicating lower values of $\mu Td$ result in better randomness before post-processing. Along with pass and fail, Dieharder tests also mention suspected results as weak. They are also included in this figure.
We have generated random numbers simulating the probability distribution given by Eq. (31) for different $\mu Td$. Three files of size $1$ GB have been generated for every $\mu Td$. Then we perform the NIST tests over the generated numbers. $100$ bit streams of length $10,000,000$ have been considered for the tests. In Fig. 15, we plot the bar charts of the $p$-values from different NIST tests. The tests for which the $p$-values are found as 00 are not included in the plot. From these plots, we can see that, as $\mu Td$ increases, the number of passed tests decreases indicating a decrement of the randomness in the random number. The overall result is plotted in Fig. 16 as the number of tests our simulated data passed. These graphs indicate that lower $\mu Td$ values give rise to better random numbers requiring less post-processing.
Figure 18: Ent test result for the data simulated with the distribution given by Eq. (31). The relative deviation is calculated as $\frac{|v_{i}-v_{o}|}{v_{i}}$, where $v_{i}$ is the expected ideal value for uniform distribution and $v_{o}$ is the observed value for the simulated data. Although the deviation of the Monte Carlo value of $\pi$, the arithmetic mean and the $\chi^{2}$ test value is high, the deviation of the entropy and the serial correlation is very small.
We have also performed the Dieharder, AIS-$31$ and ENT over the generated numbers. Figure 17 shows how many tests for our simulated data passed for Dieharder and AIS tests. Similar to the NIST tests, these tests also show that small $\mu Td$ values provide better randomness before post-processing. ENT test does not mention pass or fail as a result. However, it provides ideal and actual values for the performed tests. In Fig. 18, we plot the relative deviation of the actual values from their ideal values for different ENT tests. In this case also, the deviations increase with the increase of $\mu Td$.
X Discussion
In this paper, we consider two perspectives for analyzing QRNG. From the manufacturer’s point of view, we discuss how to determine when the source can be considered as quantum. Here we assume that photon detection time is available as a sample. We suggest two-fold statistical tests. The first one may be an interval estimation or hypothesis test to decide between sub-Poissonian and super-Poissonian. On the other hand, the last one, which is a $\chi^{2}$-test ensures Poisson distribution. This test can be performed for any quantum-photonic device to check its quantumness.
Then we discuss photon statistics for time-bin-based QRNG in practical scenarios. We consider two photon-arrival-time-based QRNG models, one using external time reference and another without such external reference. We show that, although for ideal devices these two models follow different distributions (one of which is perfectly random), these two QRNGs follow a similar distribution when the photon source is not an actual single photon source. We derive a relation when the photon statistics becomes $\epsilon$-random. We found that by making the product of the expected number of photons ($\mu T$) within an external reference cycle ($T$) and the detection efficiency ($d$) small enough, the randomness of the generated raw random number can be increased. Therefore it would require less post-processing, and it would retain maximum randomness from the quantum source. This discussion would help QRNG manufacturers to design better QRNG.
Although small $\mu Td$ provides better randomness, from the user’s perspective it may not be suitable. The random number generation rate, the timing error in registering photon detection time and the cost of the QRNG restrict us from choosing small values for $\mu,T$ or $d$. A lower expected photon count improves quantumness but comes with increased costs and a reduced generation rate. On the other hand, a shorter external reference cycle enhances quantumness but must surpass a minimum threshold to reduce timing errors, having only slight effects on cost and rate. Reduced detection efficiency boosts quantumness and lowers costs, but also decreases the generation rate. The trade-off among the randomness sourced from quantum systems, the rate of random number generation, and the cost of the QRNG is contingent upon the specific requirements of the users.
Finally, we simulated random numbers according to the photon statistics we derived and applied statistical tests including NIST, Dieharder, AIS-31, and ENT to these numbers. The results from these tests validate our theoretical conclusions based on $\mu Td$ about the randomness of the quantum sources.
In practical cases, the dark count of a detector also contributes to the generation of random numbers. In our discussion, we did not consider the dark count. This work can be extended to see the effect of dark counts on the photon statistics. Here, we have considered two models of photon arrival-time-based QRNGs, which fall in the category of trusted devices. A similar analysis can be performed for other QRNG models.
One may note that our discussions are not meant for black-box testing for commercially available QRNGs. Rather, the physicists and engineers may use these in the QRNG design phase before post-processing, for evaluation, fine-tuning and performance optimization.
Appendix A A Comparison between trusted QRNG and DI-QRNG
A device-independent quantum protocol uses quantum entanglement to ensure security. This can be verified using Bell inequality, quantum discord, etc. acin2016certified ; Alicki_2008 ; Alicki_2008a ; Martin_2023 . On the other hand, if a protocol relies on single photons as a resource, the absence of entanglement prevents the application of the validation procedures mentioned above. However, we can still verify its quantumness using photon-count statistics PhysRevA.41.475 ; fox2006quantum ; RevModPhys.68.127 ; qute.202100062 , discussed in Section  II and B, to improve the security. A detailed comparison indicating advantages and disadvantages is given in Table 2.
Trusted QRNG Device-independent QRNG
Advantages of DI-QRNG 1. Required trusted assumptions. 1. Trusted assumptions are not required or partially required for semi-DI-QRNG.
2. Security relies on trusted assumptions. Hence, less secure. 2. Security relies on Bell violation or similar quantum principles acin2016certified ; Alicki_2008 ; Alicki_2008a ; Martin_2023 . Hence, more secure.
Advantages of Trusted QRNG 1. No entanglement required. Hence it is cheaper. 1. Necessity of entanglement makes it costly.
2. No Bell test required. 2. Loophole-free Bell test is needed and performing a loophole-free Bell test is difficult PhysRevA.47.R747 ; GISIN1999103 ; PhysRevA.66.042111 ; PhysRevA.81.032106 ; Larsson_2014 .
3. Can be produced in the very compact form of USB or chip. 3. To avoid locality loophole in the Bell test, two entangled particles must be space-like separated PhysRevLett.115.250402 ; storz2023loophole . This forces the device to be either large or trusted during the Bell test.
4. Compatible for portable devices. 4. Incompatible in use cases with portable devices, e.g., automobiles, satellites, point of sale (POS) terminals, portable phones, laptops, etc. quantis
Table 2: A comparison between trusted and Device-independent QRNGs indicating their advantages and disadvantages.
From Table 2, one can easily conclude that although DI-QRNG provides more security, its size is not scalable to integrate it in a small space. Thus it is not suitable for many application scenarios requiring portable devices. Also, the high cost of DI-QRNG compared to trusted QRNG makes it incompatible with some uses like lottery games, token generations, etc.
Appendix B Relationship between Non-classicality of Light, antibunching of Photons and sub-Poissonian Distributions
The second-order correlation function fox2006quantum of light is defined by
$g^{2}(\tau):=\frac{\langle I(t)I(t+\tau)\rangle}{\langle I(t)\rangle\langle I(t+\tau)\rangle},$ (55)
[…]
$g^{2}(0)$ $=\frac{\langle I(t)^{2}\rangle}{\langle I(t)\rangle^{2}}\geq 1,$ (56)
and $\text{and }g^{2}(\tau)$ $=1\text{ for }\tau\gg\tau_{c},$
$\tau_{c}$ being the coherence time of the source.
Now, consider that a photon stream is sent through a beam splitter and two detectors $D_{0}$ and $D_{1}$ are kept in two output paths. Than the correlation function (55) can be written as
$g^{2}(\tau):=\frac{\langle n_{1}(t)n_{2}(t+\tau)\rangle}{\langle n_{1}(t)\rangle\langle n_{2}(t+\tau)\rangle},$ (57)
where $n_{i}(t)$ is the number of counts registered on detector $i$ at time $t$. If the incoming light consists of a photon stream with long intervals between two consecutive photons, the product $n_{1}(t)n_{2}(t)$ is always 00, giving $g^{2}(0)=0<1$, which is not possible with classical lights. Thus, depending on $g^{2}(t)$, light is distinguished in the following three classes fox2006quantum :
  1. (i)
    bunched light, when $g^{2}(0)>1$,
  2. (ii)
    coherent light, when $g^{2}(0)=1$,
  3. (iii)
    antibunched light, when $g^{2}(0)<1$.
Out of these three, antibunched light is undoubtedly non-classical.
The relation between photon count distribution and $g^{2}(\tau)$ is given by PhysRevA.41.475
[…]
In the above equation, if $g^{2}(\tau)\leq 1$ for all $\tau$, it implies $\langle(\Delta n)^{2}\rangle<\langle N\rangle$, i.e., a sub-Poissonian statistics, indicating non-classicality of light. However, if $g^{2}(\tau)\leq 1$ for some $\tau$, there is no direct relation between the sub-Poissonian photon count statistic and the antibunching of photons. Xou and Mandel PhysRevA.41.475 showed that if a source emits photons in two different nonvacuum modes, both having occupation number $n/2$, but different frequencies $\omega_{1}$ and $\omega_{2}$, the second-order correlation function takes the form
$g^{2}(\tau)=\frac{1}{2}\cos(\omega_{1}-\omega_{2})\tau-\frac{1}{n}+1,$ (59)
for $\tau<T$. Also, if $P(N,t,t+T)$ be the distribution of photon count in the time interval from $t$ to $t+T$,
$\langle(\Delta N)^{2}\rangle-\langle N\rangle=\langle N\rangle^{2}\left[\left(\frac{\sin(\omega_{1}-\omega_{2})T/2}{(\omega_{1}-\omega_{2})T/2}\right)^{2}-\frac{1}{n}\right],$ (60)
where $\langle N\rangle$ and $\langle(\Delta N)^{2}\rangle$ are respectively the mean and the variance of the above distribution. Choosing $T=\frac{2\pi}{\omega_{1}-\omega_{2}}$, it can be seen that $\langle(\Delta N)^{2}\rangle<\langle N\rangle$, i.e., the light is sub-Poissonian although $g^{2}(0)=(\frac{3}{2}-\frac{1}{n})>1$ for $n>2$ denoting the light is bunched. This example also shows that although both the antibunching of photons and the sub-Poissonian distribution of photon count indicate the non-classicality of light, they are not identical.
Appendix C Statistical Tools
C.1 Interval Estimation
In statistics, an interval $(a,b)$ is called the confidence interval for a distribution parameter $\Theta$, if $P(A<\Theta<B)=1-\epsilon$ for some predefined $\epsilon\in(0,1)$, where $A,B$ are the random variables corresponding to the statistics $a,b$ (they are functions of sample values) respectively. The quantity $1-\epsilon$ is called confidence level of the interval. The procedure of finding the confidence interval for the parameter $\Theta$ is called interval estimation. A statistic $a$ is called an unbiased estimate of an unknown parameter $\Theta$ if the expectation $\langle A\rangle=\Theta$. A statistic $a$ is called a consistent estimate of a parameter $\Theta$ if for every real $\epsilon>0$, $\lim\limits_{n\to\infty}P(|A-\Theta|\geq\epsilon)=0$, where $n$ is the sample size.
C.2 Hypothesis Testing
In statistics, a hypothesis is a statement or an assumption about a population distribution. this may be right or wrong. Hypothesis testing is a procedure to decide whether a statistical hypothesis, called the null hypothesis $H_{0}$, should be rejected or not. A critical region or rejection region of hypothesis testing is a specific region of test statistic values, for which the null hypothesis is rejected. The test statistic for a hypothesis test is a function of the testing sample casella2002statistical . A hypothesis test is called one-tailed if the critical region appears only at one tail of the reference distribution of the test statistic. On the other hand, it is called a two-tailed test, if the critical regions appear at both tails of the above distribution.
During hypothesis testing, two types of error may occur. Type I error occurs when $H_{0}$ is rejected given that it is true. Type II error occurs when $H_{0}$ is accepted given that it is false. Let $\alpha$ and $\beta$ be the probabilities of Type I and Type II errors respectively. The value $\alpha$ denotes the area of the critical region and is also called the significance level. The value $1-\beta$, that is, the probability that $H_{0}$ is rejected given that it is false, is also called the power of a test.
A hypothesis test starts with two hypotheses: the null hypothesis ($H_{0}$) and the alternative hypothesis ($H_{a}$). Depending on the significance level $\alpha$, the critical region is decided. The acceptance or rejection decision of $H_{0}$ is based on what is called the $p$-value of the test statistic. It is defined as $P(|X|>|x|)$, where $X$ is a random variable, whose distribution is the reference distribution of the test statistic of a hypothesis test, and $x$ is the observed value of this statistic ross2009introduction ; hogg2019introduction . When the observed test statistic falls within the critical region, $p$-value of the test statistic becomes less than $\alpha$. In this case, $H_{0}$ is rejected. If the $p$-value of the test statistic is greater than $\alpha$, the observed test statistic falls outside the critical region. In that case, $H_{0}$ will be accepted.
C.3 Goodness-of-fit Test
Goodness-of-fit test is a statistical procedure to decide whether sample data comes from a particular distribution or not. This is a specific hypothesis testing where
$H_{0}$: Sample data comes from a distribution to be tested.
$H_{a}$: Sample data not from that distribution.
Chi-square ($\chi^{2}$) test10.1214/aoms/1177729380 , Kolmogorov–Smirnov test (KS test) 16e7f618-c06b-3d10-8705-1086b218d827 , and Anderson–Darling test 10.1214/aoms/1177729437 are some examples of the goodness-of-fit test good . Out of these tests, the last two tests are designed for continuous distributions only. However, the KS test is modified later to consider discrete distributions as well 45e251e1-e12b-3bce-b1a2-da8f4c1c2382 . The $\chi^{2}$ test has been found more robust than the KS test for the continuous distributions 32ba378b-c6d3-36dc-9d3e-0b9814b089a4 .
Appendix D Quantum Entropy for QRNG
The entropy generated by an RNG is the source of randomness for its outcome. The outcomes of an RNG follow a probability distribution, say $P(X)$, where $X$ is the underlying random variable.
Definition 3.
The Shannon entropy 6773024 ; 6773067 of a discrete random variable $X$ is defined as
$H(X)=-\sum_{x}P(X=x)\log P(X=x).$ (61)
The minimum entropy of an RNG is the smallest amount of the entropy that the RNG can produce.
Definition 4.
Minimum entropy $H_{min}$ of raw data generated by a RNG is defined as
$H_{min}:=-\log\left(\max_{x}P(X=x)\right),$ (62)
where $X$ is the underlying discrete random variable.
This minimum entropy is used to decide the amount of post-processing required for transforming the raw data generated by an RNG to uniform random numbers.
The Shannon entropy of the probability distribution in Eq. (31) is given by
$H(Q)=$ $-\sum_{i=1}^{N}f(i)\log f(i)$
$=$ $-\sum_{i=1}^{N}\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}e^{-i\frac{\mu Td}{N}}\log\left(\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}e^{-i\frac{\mu Td}{N}}\right)$
$=$ $-\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}\sum_{i=1}^{N}e^{-i\frac{\mu Td}{N}}\left[\log\left(e^{-i\frac{\mu Td}{N}}\right)+\log\left(\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}\right)\right]$
$=$ $\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}\left[\log\left(e^{\frac{\mu Td}{N}}\right)\sum_{i=1}^{N}ie^{-i\frac{\mu Td}{N}}-\log\left(\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}\right)\sum_{i=1}^{N}e^{-i\frac{\mu Td}{N}}\right]$
$=$ $\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}\Bigg{[}\log\left(e^{\frac{\mu Td}{N}}\right)\frac{e^{-\frac{\mu Td}{N}}\left[1-(N+1)e^{-\mu Td}+Ne^{-(N+1)\frac{\mu Td}{N}}\right]}{(1-e^{-\frac{\mu Td}{N}})^{2}}$
$\;-\log\left(\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}\right)\frac{e^{-\frac{\mu Td}{N}}(1-e^{-\mu Td})}{1-e^{-\frac{\mu Td}{N}}}\Bigg{]}$
$=$ $\frac{1-(N+1)e^{-\mu Td}+Ne^{-(N+1)\frac{\mu Td}{N}}}{(1-e^{-\mu Td})(1-e^{-\frac{\mu Td}{N}})}\log\left(e^{\frac{\mu Td}{N}}\right)-\log\left(\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}\right).$ (63)
This is the entropy generated by the photon arrival-time-based QRNGs we have discussed in this paper.
Appendix E Randomness Testing Suites
Randomness testing is very important to ensure whether an RNG is producing perfect random numbers or not. It tells us about the uniformity of the underlying distribution that the outputs of an RNG follow.
Definition 5.
Let $S_{k}=\{0,1\}^{k}$ be the set of all $k$-bit binary strings containing $2^{k}$ elements. A $n$-bit string, $X=x_{1}x_{2}\dots x_{n},\ x_{i}\in\{0,1\}$ for all $i$, is called perfect (or uniformly) random if and only if for any $k$-bit ($k\leq n$) sub-string $x=x_{j}x_{j+1}\dots x_{j+k}$, for some $j\leq n-k$, of this string, the probability that $x=y$ for any $y\in S_{k}$ is $\frac{1}{2^{k}}$. That is,
$X=x_{1}x_{2}\dots x_{n},x_{i}\in\{0,1\}\ \forall i,\text{ is true random}$
$\iff$ $\forall k<n,\forall j\leq n-k\text{ and }\forall y\in S_{k},\;P(x=y)=\frac{1}{2^{k}},$
where $x=x_{j}x_{j+1}\dots x_{j+k}$.
This definition of a perfect random number uses probability over all possible combinations of the bit string to ensure randomness. However, in a practical case, when we have only a finite sample, this probability cannot be perfectly calculated. Hence this definition cannot be directly used to test the randomness of random numbers.
To overcome this limitation, several randomness testing suites like NIST 10.5555/2206233 , Dieharder Die , AIS-$31$ killmann2011proposal , ENT ENT etc. have been proposed. Recently, Foreman et al. also suggested some test suite foreman2024 for testing randomness. Each of these suits consists of multiple tests to decide the randomness of the input data. For each test, they calculate the value of the test statistic, and depending on these values the randomness is decided. NIST and Dieharder perform hypothesis testing with the null hypothesis that the number is random, AIS-31 performs interval estimation, and ENT directly returns the test statistic values. These suites consider different aspects, such as entropy, serial correlation, frequency of values, repetition of similar patterns etc. to perform the test. Some tests also perform Monte Carlo simulation MonteCarlo to validate the randomness.